### v1

# Approximate Degrees of Multisymmetric Properties with Application to Quantum Claw Detection

Abstract
The claw problem is central in the fields of theoretical computer science as well as cryptography. The optimal quantum query complexity of the problem is known to be $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ for input functions $f\colon[F]\to Z$ and $g\colon[G]\to Z$. However, the lower bound was proved when the range $Z$ is sufficiently large (i.e., $\lvert Z\rvert=\Omega(FG)$). The current paper proves the lower bound holds even for every smaller range $Z$ with $\lvert Z\rvert\geq F+G$. This implies that $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for every such range. In addition, the lower bound $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ is provided for even smaller range $Z=[M]$ with every $M\in[2,F+G]$ by reducing the claw problem for $\lvert Z\rvert=F+G$. The proof technique is general enough to apply to any $k$-symmetric property (e.g., the $k$-claw problem), i.e., the Boolean function $\Phi$ on the set of $k$ functions with different-size domains and a common range such that $\Phi$ is invariant under the permutations over each domain and the permutations over the range. More concretely, it generalizes Ambainis’s argument [Theory of Computing, 1(1):37-46] to the multiple-function case by using the notion of multisymmetric polynomials.
1 Introduction
1.1 Background
Due to their simplicity and broad applicability, the collision detection/finding problem and its variant, the claw detection/finding problem, are central problems in the fields of theoretical computer science (e.g.,[14, 20, 2, 6, 10, 31]) as well as cryptography (e.g.,[13, 23, 27, 25, 17]). For given function $f$, the goal of the collision detection is to detect the existence of a pair of elements $(x,y)$ such that $f(x)=f(y)$ (since the collision detection without any promise on the input $f$ is often called the element distinctness, we may use the latter term when it is appropriate). As a natural variant of the collision detection, we can define the claw detection for two functions $f,g$ as the problem of detecting the existence of a pair $(x,y)$ such that $f(x)=g(y)$. When we are required to find a collision (a claw) if it exists, we refer to the corresponding “search” problem as the collision finding (the claw finding, respectively) in this paper.
For large domain(s), these problems are generally hard (i.e., time-consuming) to solve. This hardness forms the basis of the security of various cryptographic applications. This fact in turn makes these problems appealing as targets with which to investigate quantum computational power: If one finds quantum algorithms that solve the problems very efficiently, this exhibits solid quantum advantages concerning practically essential problems; if one proves the hardness of the problems even against quantum computation, this gives theoretical evidence that the cryptographic applications are secure against quantum computers.
Interestingly, the actual situation is somewhere in between. For these problems, quantum speed-ups are possible but at most polynomials. This has motivated the invention of various important techniques for deriving quantum upper bounds (i.e., quantum algorithms) and lower bounds.
For the element distinctness, Buhrman et al. [14] proved the upper bound $O(n^{3/4})$ and the lower bound $\Omega(\sqrt{n})$ on the quantum query complexity for $f\colon[n]\to Z$ for a finite set $Z$. Aaronson and Shi [2] improved the lower bound to $\Omega(n^{2/3})$ by introducing a novel technique to the polynomial method to bound the degrees of multi-variate polynomials that approximate symmetric Boolean functions, where $[n]$ denotes the set $\{1,\dots,n\}$. Ambainis [6] then provided the breakthrough upper bound $O(n^{2/3})$ by developing a general framework of quantum search based on quantum walk. His work was followed by many generalizations and variants that form one of the mainstreams of quantum algorithms (e.g.,[29, 26, 8, 22]).
As for the claw detection for $f\colon[F]\to Z$ and $g\colon[G]\to Z$, where $F$ and $G$ are positive integers such that $F\leq G$ and $Z$ is a finite set, Buhrman et al. [14] proved that the quantum query complexity is $\Omega(\sqrt{G})$ and $O(\sqrt{G}+F^{1/2}G^{1/4})$. Zhang [32] then proved the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ on the quantum query complexity by providing the reduction from the collision detection with promised that the input function with domain $[FG]$ is either one-to-one or two-to-one and using the lower bound $\Omega((FG)^{1/3})$ [2] for that promised case. Tani [30] then proved the matching upper bound $O(\sqrt{G}+(FG)^{1/3})$ by running a quantum-walk-based search on a particular graph product of two Johnson graphs corresponding to the two input functions.
These tight bounds depend only on the domain size(s) of the input function(s). This means that these bounds are tight if the function range is sufficiently large. Indeed, the lower bound $\Omega(n^{2/3})$ of the element distinctness was originally proved when the range size is at least $1.5n$ [2]. In some cases of smaller range size, however, it is still non-trivial to detect a collision (e.g., a collision for a given function $f\colon[n]\to[n]$ since $f$ can be one-to-one). Moreover, in some applications, such as cryptographic ones, the range size is often known and can be a relatively small value. Ambainis [4] and Kutin [24] independently proved the same lower bound $\Omega(n^{2/3})$ holds even for the range of size $n$. In particular, Ambainis’s argument is so general that it can be applied to any function $\Phi$ from a set of functions $f$ to $\{0,1\}$ if $\Phi$ is invariant under the permutations $\pi$ and $\sigma$ over the domain and range, respectively, of $f$, i.e., $\Phi(f)=\Phi(\sigma f\pi)$ (e.g., if $\Phi$ is the function that tells the existence of a collision for $f$, then it is easy to see that $\Phi(f)=\Phi(\sigma f\pi)$).
As for the claw detection, Zhang’s lower bound holds if the range size of input functions is $\Omega(FG)$. This condition seems very strong, considering that it is still non-trivial to detect a claw for a given pair of functions $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G$, since the images of $f$ and $g$ can be disjoint. Therefore, it is natural to ask whether the lower bound $\Omega(\sqrt{G}+(FG)^{1/3})$ still holds for a smaller range, such as $[F+G]$, and, more ambitiously, whether a multi-function version of Ambainis’s argument exists. From a practical point of view, relaxing the condition on the function range for tight bounds could contribute to more accurate security analysis of post-quantum ciphers, such as those based on supersingular elliptic curve isogenies [21], one form (SIKE) of which had been considered until the final round of the post-quantum standardization at NIST.
1.2 Polynomial Method
There are two primary lower-bound techniques for quantum query complexity: the polynomial method [9] and the adversary method [3, 18]. Although the polynomial method does not necessarily give optimal lower bounds [5], it often gives non-trivial lower bounds that are not known how to obtain using other lower-bound techniques. The highlight of the method is the claim that the minimum degree of a polynomial that approximates a Boolean function $\Phi$ is at most twice the number of quantum queries required to compute $\Phi$. This allows us to use the results of polynomial approximation theory to obtain lower-bounds on the quantum query complexity of $\Phi$. More concretely, let $f$ be any function from $[n]$ to $[m]$. Then, $f$ can be identified with $nm$ Boolean variables $\{x_{ij}:i\in\left[n\right],j\in\left[m\right]\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$. Let $\Phi$ be a Boolean function that tells whether input $f$ has a certain property (e.g., the property that $f$ has at least one collision). Then, $\Phi$ is a function in the variables $\{x_{ij}\}$. The polynomial method claims that the minimum degree of a polynomial that approximates $\Phi$ within error $\epsilon$ is at most twice the number of quantum queries required to compute $\Phi$ with error probability at most $\epsilon$. For $m\geq n$, suppose that $\Phi^{\prime}$ is the restriction of $\Phi$ to the case where the range of input function is $\left[n\right]$. Then, it is easy to see that if $d$ is the minimum degree of a polynomial that approximates $\Phi$, then the minimum degree of a polynomial that approximates $\Phi^{\prime}$ is at most $d$ (since $\Phi^{\prime}$ is obtained by setting $x_{ij}=0$ for all $i\in[n]$ and $j\in\left[m\right]\setminus\left[n\right]$ in $\Phi$). Ambainis [4] proved that the converse is also true if $\Phi$ is a symmetric property, i.e., a Boolean function invariant under the permutations over domain $[n]$ and range $\left[m\right]$, respectively, of $f$. For the property $\Phi$ that input $f\colon[n]\to[3n/2]$ has at least one collision, the minimum approximate degree of $\Phi$ is known to be $\Omega\left(n^{2/3}\right)$. Thus, the polynomial method together with Ambainis’s argument implies that the quantum query complexity of the element distinctness problem is $\Omega\left(n^{2/3}\right)$ even if the function range is $\left[n\right]$. This relaxes the condition (i.e., $m\geq 1.5n$ ) on the range size of input functions required in [2] to obtain the optimal lower bound.
1.3 Our Contribution
Our main contribution is to prove a multi-function version of Ambainis’s argument. As an application of our argument to two functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, we show that the optimal quantum lower bound for claw detection proved for range size $FG$ still holds for smaller range size $F+G$. Moreover, by using this new lower bound, we obtain a lower bound for an even smaller range $[M]$ with $2\leq\left|M\right|<F+G$. In the following, we summarize our results more concretely. To simplify the notation, we represent by $\mathcal{R}^{\mathcal{D}}$ the set of functions from $\mathcal{D}$ to $\mathcal{R}$ for any ordered sets $\mathcal{D}$ and $\mathcal{R}$. For instance, if $f$ is a function from $[F]$ to $[M]$, then $f\in[M]^{[F]}$.
1.3.1 Approximate Degree of Multisymmetric Properties
For simplicity, we first consider the case where two functions $f\colon\left[F\right]\rightarrow\left[M\right]$ and $g\colon\left[G\right]\rightarrow\left[M\right]$ are given as input. Then, $f$ can be identified with $FM$ Boolean variables $\{x_{ij}:i\in\left[F\right],j\in\left[M\right]\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$. Similarly, $g$ can be identified with $GM$ Boolean variables $\{y_{kj}:k\in\left[F\right],j\in\left[M\right]\}$. For $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, Let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a Boolean function that tells whether input pair $(f,g)$ in $\mathcal{D}$ has a certain property (e.g., the claw property that $f$ and $g$ have at least one claw). Then, $\Phi$ is a function in the $FM+GM$ variables $\{x_{ij}\}\cup\{y_{kj}\}$. We say that $\Phi$ is bisymmetric, if $\Phi$ is invariant under the permutations over domains $\left[F\right]$ and $[G]$, respectively, and the common range $\left[M\right]$ of $f$ and $g$, i.e., $\Phi\left(f,g\right)=\Phi\left(\sigma f\pi_{F},\sigma g\pi_{G}\right)$, where $\pi_{F}$ and $\pi_{G}$ are permutations over $[F]$ and $[G]$, respectively, and $\sigma$ is a permutation over $\left[M\right]$. For instance, if $\Phi$ is the claw property, then $\Phi$ is bi-symmetric.
Theorem 1
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$, let $\Phi\colon[M]^{[F]}\times[M]^{[G]}\to\{0,1\}$ be a bisymmetric Boolean function of a pair of functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$. In addition, for every $M^{\prime}$ with $F+G\leq M^{\prime}<M$, $\Phi^{\prime}\colon[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\to\{0,1\}$ be the restriction of $\Phi$ to the case where the input functions are $f\colon[F]\to[M^{\prime}]$ and $g\colon[G]\to[M^{\prime}]$. Then, for $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.
With the promise that the union of the images of input functions $f$ and $g$, i.e., ${f([F])\cup g([G])}$, is upper-bounded in size by $\mathfrak{U}$, Theorem 1 also holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to the proof. For instance, if both $f$ and $g$ are two-to-one, then $\lvert f([F])\cup g([G])\rvert\leq(F+G)/2$, and thus Theorem 1 holds for every $M^{\prime}$ with $(F+G)/2\leq M^{\prime}<M$. Since $\lvert f([F])\cup g([G])\rvert\leq F+G$ always holds, we obtain Theorem 1 in the above form (i.e., without any promise on the image sizes).
The definition of bisymmetric properties can be generalized in a straightforward way: Let $\Phi\colon\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}\to\{0,1\}$ be a Boolean function over the set of $k$-tuples of functions in $\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\colon[F_{\ell}]\to[M]$ for $F_{\ell},M\in\mathbb{Z}_{+}$. As in the case of two functions, we can see that $\Phi$ depends on Boolean variables $x^{(\ell)}_{ij}\in\{0,1\}$ for $(\ell,i,j)\in[k]\times[F_{\ell}]\times[M]$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant under the permutations over $[F_{\ell}]$ for each $\ell\in[k]$, respectively, and the common range $[M]$ of $f_{\ell}$’s, i.e.,
$\Phi(f_{1},\dots,f_{k})=\Phi\left(\sigma f_{1}\pi_{F_{1}},\dots,\sigma f_{k}\pi_{F_{k}}\right),$
for every permutation $\pi_{F_{i}}$ over $[F_{i}]$ for $i\in[k]$ and every permutation $\sigma$ over $[M]$. Then, we have a generalization of Theorem 1.
Theorem 2
Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, $\Phi:\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}\to\{0,1\}$ be a $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell$. In addition, for every ${M^{\prime}}$ such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[{M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}$, where $\mathcal{F}^{\prime}_{\ell}$ denotes $[{M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$.
Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$
With the promise that the union of the images of input functions, i.e., $\bigcup_{\ell\in[k]}{f_{\ell}([F_{\ell}])}$, is upper-bounded in size by $\mathfrak{U}$, Theorem 2 also holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to the proof. Slightly general versions of the above theorems are Theorems 12 and 15 proved in Sec. missing  3.
1.3.2 Application to Claw Detection
$2\leq M<\max\{2,(G/F)^{2}\}$ $\max\{2,(G/F)^{2}\}\leq M<F+G$ $F+G\leq M<FG$ $FG\leq M$
$F\leq G\leq F^{2}$ $O\left(\min\left\{(FG)^{1/3},G^{1/2+\epsilon}M^{1/6}\right\}\right)$ [30, 7] $O((FG)^{1/3})$ [30]
$\Omega(\sqrt{G})$ [14] $\Omega(F^{1/3}G^{1/6}M^{1/6})$ [Th.5] $\Omega((FG)^{1/3})$ [Th.4] $\Omega((FG)^{1/3})$ [32]
$F^{2}<G$ $\Theta\left(\sqrt{G}\right)$ [14]
Table 1: Summary of quantum query complexity for claw detection for functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, where $F\leq G$. Note that $(G/F)^{2}<F+G$ holds if $G\leq F^{2}$ (otherwise, $\Theta(\sqrt{G})$ is tight for all $M\geq 2$).
We use Theorem 1 to derive a lower bound for the claw detection with a small range. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property (for input functions with a possibly large common range). It turns out that for any constant $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs $f\colon[F]\to[FG]$ and $g\colon[G]\to[FG]$ is $\Omega\left((FG)^{1/3}\right)$. By combining this with Theorem 1, we can obtain a lower bound on the minimum approximate degree of the claw property for the range size at least $F+G$ of input functions.
Theorem 3
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $20<\epsilon<1/$ 2 and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for a function pair in $[M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
By the polynomial method, Theorem 3 implies the optimal lower bound on the bounded-error quantum query complexity of both the claw detection and the claw finding, where the optimality follows from the quantum algorithm [30] that finds a claw.
Theorem 4
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $[M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
Ambainis, Balodis and Iraids [7] investigated upper and lower bounds for the claw problem for input function pairs with much smaller ranges. They proved the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ when the two input functions has the same domain size $N$ and the range size $M$ with $2\leq M<N$. Their proof is based on the lower bound $\Omega(N^{2/3})$ on the number of quantum queries required to detect a claw for a given pair of functions in $[N]^{[N]}\times[N]^{[N]}$, which follows by reducing from the element distinctness. However, if the domains of two input functions are different in size, their proof breaks down. By generalizing their proof using Theorem 4, we obtain a lower bound for functions with different-size domains and very small common ranges.
Theorem 5
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and $M<F+G$. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $[M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
$\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)=\left\{\begin{array}[]{lcl}\Omega(\sqrt{G})&\text{if}&2\leq M<(G/F)^{2},\\ \Omega(F^{1/3}G^{1/6}M^{1/6})&\text{if}&(G/F)^{2}\leq M<F+G.\end{array}\right.$
Note that the condition of $F\leq G\leq F^{2}$ and $M<F+G$ imposed in Theorem 5 means the parameter range that Theorem 4 does not cover. Setting $F=G=N$ in the theorem recovers the lower bound $\Omega\left(N^{1/2}M^{1/6}\right)$ [7].
To evaluate our bound in Theorem 5, we will compare it with the lower bound obtained by straightforwardly reducing the claw detection for two functions with domains of the same size and applying the lower bound given in  [7]. Let us write $\mathsf{claw}_{(a,b)\to c}$ to mean the decision problem of detecting the existence of a claw for given two functions in $[c]^{[a]}\times[c]^{[b]}$. The reduction from $\mathsf{claw}_{(G,G)\to M}$ to $\mathsf{claw}_{(F,G)\to M}$ is as follows, where we assume for simplicity that $G$ is divisible by $F$, i.e., $F|G$ (it is easy to extend the reduction to the general case by padding dummy elements to one of the domains). Let $(f,g)$ be a pair of input functions to $\mathsf{claw}_{(G,G)\to M}$. We first partition the domain $[G]$ of $f$ into $G/F$ blocks of size $F$: $B_{1},\dots,B_{G/F}$. For randomly chosen $i\in[G/F]$, we run a $Q$-query quantum algorithm for $\mathsf{claw}_{(F,G)\to M}$ on $f$ restricted to the domain $B_{i}$ and $g$, and apply the quantum amplitude amplification [12, 19] to the whole algorithm. The resulting quantum algorithm makes $O\left(\sqrt{G/F}\cdot Q\right)$ quantum queries. Since the quantum query complexity of $\mathsf{claw}_{(G,G)\to M}$ for $M<G$ is $\Omega(G^{1/2}M^{1/6})$ [7], we have $Q=\Omega(F^{1/2}M^{1/6})$. Together with trivial lower bound $\Omega(\sqrt{G})$, we obtain the lower bound $\Omega(\sqrt{G}+F^{1/2}M^{1/6})$. This bound is weaker than our $\Omega(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6})$. To see this, assume $G=F^{1+\epsilon}$ for $0<\epsilon\leq 1$, satisfying $F<G\leq F^{2}$. In this case, we have
$F^{\frac{1}{3}}G^{\frac{1}{6}}M^{\frac{1}{6}}=F^{\frac{1}{3}}F^{\frac{1}{6}(1+\epsilon)}M^{\frac{1}{6}}=F^{\frac{1}{2}+\frac{1}{6}\epsilon}M^{\frac{1}{6}}\in\omega\left(F^{\frac{1}{2}}M^{\frac{1}{6}}\right).$
For instance, if $F=G^{3/4}$ and $M=G^{3/4}$, then our bound is $\Omega(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6})=\Omega(G^{13/24})$ while the straightforwardly obtained bound is $\Omega(\sqrt{G}+F^{1/2}M^{1/6})=\Omega(\sqrt{G})$.
Tab. missing 1 summarizes the quantum query complexity for claw detection for functions $f\colon[F]\to[M]$ and $g\colon[G]\to[M]$, assuming $F\leq G$ without loss of generality. The upper bound $O(G^{1/2+\epsilon}M^{1/4})$ in the case of $M<F+G$ for $F\leq G\leq F^{2}$ follows from the upper bound [7] for $\mathsf{claw}_{(G,G)\to M}$ (i.e., the claw detection for the same domain sizes) via a simple embedding the domain $[F]$ into $[G]$.
The most interesting open problem is obviously to find the tight bound in the case of $M\leq F+G$ for $F\leq G\leq F^{2}$. More precisely, it follows from Tab. missing 1 that the tight bound $\Theta((FG)^{1/3})$ for $M\geq F+G$ also holds for all $M=\Omega(G)$. On the other hand, if $M=\Theta(1)$, then $\Omega(\sqrt{G})$ is tight, since Grover’s search algorithm achieves $O(\sqrt{G})$ queries. Thus, the tight bound is open when $M$ is asymptotically smaller than $G$ and larger than $2$.
1.4 Technical Outline
Approximate Degree of Multisymmetric Properties.
The starting point of the proofs of Theorems 1 and 2 is the proof idea for the single-symmetric properties introduced by Ambainis [4] (we hereafter refer to "single-symmetric" just as "symmetric"). Let $\Phi$ and $\Phi^{\prime}$ be the symmetric Boolean functions defined in Sec. missing  1.2. Recall that $\Phi$ and $\Phi^{\prime}$ are functions in $\{x_{ij}\colon(i,j)\in[n]\times[m]\}$ and $\{x_{ij}\colon(i,j)\in[n]\times[n]\}$, respectively. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$. Then, Ambainis’s proof consists of two claims. The first claim is that the following are equivalent:
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $z_{1},\ldots,z_{m}$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $P$ of degree at most $d$ in $x_{11},\ldots,x_{nm}$ that $\epsilon$-approximates $\Phi$.
The second claim is that the following are equivalent:
  1. i
    There exists a polynomial $Q\left(z_{1},\ldots,z_{m}\right)$ of degree at most $d$ that $\epsilon$-approximates $\Phi$.
  2. ii
    There exists a polynomial $Q^{\prime}\left(z_{1},\ldots,z_{n}\right)$ of degree at most $d$ that $\epsilon$-approximates $\Phi^{\prime}$.
These claims together imply that the minimum degree of a polynomial that approximates $\Phi$ is at most $d$, if and only if the minimum degree of a polynomial that approximates $\Phi^{\prime}$ is at most $d$.
We want to generalize these two claims to the bisymmetric case (the $k$-symmetric case is a straightforward extension of the bisymmetric case). Recall in the bisymmetric case that $f\colon[F]\to[M]$ is identified with $FM$ Boolean variables $\{x_{ij}:i\in\left[F\right],j\in\left[M\right]\}$, and $g\colon[G]\to[M]$ is identified with $GM$ Boolean variables $\{y_{kj}:k\in\left[F\right],j\in\left[M\right]\}$. In adition, we define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Let $\Phi$ and $\Phi^{\prime}$ be the bisymmetric Boolean function in $\{x_{ij}\}$ and $\{y_{kj}\}$ as defined in Sec. missing  1.3. The first claim in the bisymmetric case is that the following are equivalent:
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $z_{1},\ldots,z_{M},w_{1},\dots,w_{M}$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $P$ of degree at most $d$ in $x_{11},\ldots,x_{FM},y_{11},\dots,y_{GM}$ that $\epsilon$-approximates $\Phi$.
The second claim in the symmetric case is that the following are equivalent for every $M^{\prime}\in[F+G,M]$:
  1. i
    There exists a polynomial $Q\left(z_{1},\ldots,z_{M},w_{1},\dots,w_{M}\right)$ of degree at most $d$ that $\epsilon$-approximates $\Phi$.
  2. ii
    There exists a polynomial $Q^{\prime}\left(z_{1},\ldots,z_{M^{\prime}},w_{1},\dots,w_{M^{\prime}}\right)$ of degree at most $d$ that $\epsilon$-approximates $\Phi^{\prime}$.
For the first claim in the symmetric case, it is easy to show that item $1$ implies item $2$ by using the trivial relation $z_{j}=\sum_{i}x_{ij}$. This works also in the bisymmetric case. To show the converse in the symmetric case, we consider the average $\mathsf{E}_{\pi}[P]$ of $P$ over all permutations $\pi$ over the domains of $f$, since $\mathsf{E}_{\pi}[P]$ also approximates $\Phi$ due to the symmetry of $\Phi$ over the domains. Since we are interested in the degree of the resulting polynomial, it suffices to examine the average of each product term $x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}$ in $P$ and it turns out that $\mathsf{E}_{\pi}\left[x_{i_{1}j_{1}}\cdots x_{i_{k}j_{k}}\right]$ is a polynomial in $z_{j}$’s of degree at most $d$. In the bisymmetric case, we take average $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$ of $P$ over all permutations $\pi_{F}$ over the domains of $f$ and all permutations $\pi_{G}$ over the domains of $g$. Note that each product term in $P$ is of the form of $x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}$. To obtain the average of each term, we apply permutations $\pi_{F}$ and $\pi_{G}$ independently and thus it holds that
$\mathsf{E}_{\pi_{F}\times\pi_{G}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right]=\mathsf{E}_{\pi_{F}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}\right]\cdot\mathsf{E}_{\pi_{G}}\left[y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right].$
It turns out $\mathsf{E}_{\pi_{F}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}\right]$ and $\mathsf{E}_{\pi_{G}}\left[y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right]$ are polynomials in $z_{j}^{\prime}s$ and $w_{j}$’s, respectively, of degree $d$ by a similar analysis to the symmetric case. Consequently, we can obtain the bisymmetric version of the first claim.
For the second claim, it is more involved to generalize the symmetric case to the bisymmetric case since the permutation over the range is common to the two functions. It is easy to see that item (i) implies (ii) in the symmetric case since we can obtain $Q^{\prime}$ by setting $z_{n+1}=\cdots=z_{m}=0$ in $Q$. This works also in the bisymmetric case. For the converse, let us first review the symmetric case. Since every $z:=(z_{1},\dots,z_{m})$ corresponding to a function $f:\left[n\right]\rightarrow\left[m\right]$ has at most $n$ nonzero coordinates, there exists a permutation $\sigma$ over $[m]$ that moves all non-zero coordinates to the first $n$ coordinates, that is, $\sigma z=\left(z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(m\right)}\right)$ satisfies that $z_{\sigma^{-1}\left(j\right)}=0$ for every $\sigma^{-1}\left(j\right)\geq n+1$. This action of $\sigma$ does not change the value of $\Phi$: $\Phi\left(f\right)=\Phi\left(\sigma f\right)$. Since $\sigma f$ can be regarded as a function in $\left[n\right]^{\left[n\right]}$, it must hold that $\Phi\left(\sigma f\right)=\Phi^{\prime}\left(\sigma f\right)$. By the assumption, $\Phi^{\prime}\left(\sigma f\right)$ is approximated by $Q^{\prime}\left(\sigma z\mid_{n}\right)$, where $\sigma z\mid_{n}$ is the first $n$ variables of $\sigma z$, i.e., $z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(n\right)}$. We then define from $Q^{\prime}$ a symmetric polynomial $Q$ in $m$ variables, $z_{1},\ldots,z_{m}$, such that, for $z$ corresponding to an arbitrary $f$,
$Q\left(z\right)=Q\left(\sigma z\right)=Q\left(z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(n\right)},0,\ldots,0\right)=Q^{\prime}\left(z_{\sigma^{-1}\left(1\right)},\ldots,z_{\sigma^{-1}\left(n\right)}\right)=Q^{\prime}\left(\sigma z|_{n}\right),$
where $\sigma$ is a particular permutation defined above depending on $f$. This implies that $Q\left(z\right)$ approxiamtes $\Phi\left(f\right)$, since $Q^{\prime}\left(\sigma z\mid_{n}\right)$ approximates $\Phi^{\prime}\left(\sigma f\right)=\Phi\left(\sigma f\right)=\Phi\left(f\right)$. Since we can assume without loss of generality that $Q^{\prime}$ is a symmetric polynomial, the fundamental theorem of symmetric polynomials implies that $Q^{\prime}$ is represented as a polynomial in the elementary polynomials over $z_{1},\dots,z_{n}$. We then replace these elementary polynomials in $Q^{\prime}$ with the corresponding polynomials over $z_{1},\dots,z_{m}$ to obtain $Q$.
In the bisymmetric case, $Q$ is a polynomial in two sets of variables: $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$; $Q^{\prime}$ is a polynomial in two sets of variables $z_{1},\dots,z_{M^{\prime}}$ and $w_{1},\dots,w_{M^{\prime}}$ such that $F+G\leq M^{\prime}<M$, since there exists a permutation $\sigma$ over $[M]$ that moves all non-zero coordinates in $z_{1},\dots,z_{M}$ and $w_{1},\dots,w_{M}$ to the first $M^{\prime}$ coordinates, respectively. To obtain $Q$ from $Q^{\prime}$, we utilize a certain symmetry of $Q^{\prime}$. However, $Q^{\prime}$ is not invariant under the action of permutations over $[2M^{\prime}]$ to $2M^{\prime}$ variables $z_{1},\dots,z_{M^{\prime}},w_{1},\dots,w_{M^{\prime}}$, but it is invariant under the action of $\tau\times\tau$ for permutation $\tau$ over $[M^{\prime}]$ such that $(z_{1},\dots,z_{M})\mapsto\left(z_{\tau^{-1}\left(1\right)},\ldots,z_{\tau^{-1}\left(M^{\prime}\right)}\right)$ and $(w_{1},\dots,w_{M})\mapsto\left(w_{\tau^{-1}\left(1\right)},\ldots,w_{\tau^{-1}\left(M^{\prime}\right)}\right)$. Here, the notion of multisymmetric polynomials comes in. Roughly speaking, multisymmetric polynomials are those invariant under the action of permutations to the set of vectors of variables, instead of variables. In our case, let us consider the set of vectors $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$, each of which consists of two variables. Then, $Q^{\prime}$ is invariant under the action of permutation $\tau$ to this set of vectors. Now that we enter the world of multisymmetric polynomials, we can use the fundamental theorem of those polynomials. More concretely, $Q^{\prime}$ is represented as a polynomial in power-sum polynomials (defined in Sec. missing  2.4) in $(z_{1},w_{1}),\dots,(z_{M^{\prime}},w_{M^{\prime}})$. Then, we can obtain $Q$ by replacing each occurrence of power-sum polynomials in $Q^{\prime}$ to the corresponding polynomials in $(z_{1},w_{1}),\dots,(z_{M},w_{M})$. This approach is easily generalized to the $k$-symmetric case. We just consider the symmetry under the action of permutations to the set of vectors $\left(z_{1}^{(1)},\dots,z_{1}^{(k)}\right),\dots,\left(z_{M^{\prime}}^{(1)},\dots,z_{M^{\prime}}^{(k)}\right)$, where $z_{i}^{(\ell)}=\left\lvert f_{\ell}(j)\right\rvert$.
Application to Claw Problem
To provide a lower bound for the claw problem with a small range using Theorem 1, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. For this, we use the Zhang’s reduction [32] from the collision problem to the claw problem. Suppose that there exists a polynomial $P^{\prime}$ of degree $d$ that 0.1-approximates the claw property for functions $f\in[FG]^{[F]}$ and $g\in[FG]^{[G]}$ such that $F\leq G$. Based on $P^{\prime}$, we will then construct a polynomial $P$ of degree $d$ that (41/86)-approximates the collision property for functions $h\in[M]^{[M]}$ for $M=FG$ with the promise that $h$ is either one-to-one or two-to-one. Here we mean by the collision property the Boolean function $\Phi$ which is $\mathsf{true}$ if and only if the input function $h$ is two-to-one. Since the minimum degree of the latter polynomial is $\Omega(M^{1/3})$ [2, 4], we obtain $d=\Omega((FG)^{1/3})$. Theorem 1 with this bound implies Theorem 3, from which we obtain Theorem 4 with the polynomial method.
To obtain polynomial $P$ from $P^{\prime}$, let $P^{\prime}_{ST}$ for disjoint subsets $S,T$ of domain $[M]$ of $h$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $[M]^{S}\times[M]^{T}$ to $\{0,1\}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T$ of size $F$ and $G$, respectively:
$P:=\mathsf{E}_{S,T\subset[M]\colon S\cap T=\emptyset}[P^{\prime}_{ST}].$
Thus, $P$ is a polynomial of degree at most $d$ in variables $x_{ij}$ for $i,j\in[M]$.
If $h$ is one-to-one, then $h|_{S}$ and $h|_{T}$ obviously have no collision for every $S$ and $T$, where $h|_{S}$ and $h|_{T}$ are the restriction of $h$ to the the subsets $S$ and $T$, respectively, of the domain. Thus, $P$ is close to 00. If $h$ is two-to-one, $h|_{S}$ is injective with high probability and, in this case, $T$ intersects the preimage of $h|_{S}(S)$ with high probability. This follows from the birthday paradox, since $|S|=F<<\sqrt{M}=\sqrt{FG}$ and $|T|=G>>\sqrt{M}$. Therefore, $P$ is far from 00. By scaling and shifting $P$, we have a polynomial that $O(1)$-approximates the collision property.
As an application of Theorem 4, we obtain Theorem 5 by generalizing the reduction from the claw problem for two functions with the same domain size [7] to the reduction from the problem for two functions with the different domain size: $f\colon[F]\to[F+G]$ and $g\colon[G]\to[F+G]$ with $F\leq G\leq F^{2}$ (the case of $G>F^{2}$ is already covered by Theorem 4). More precisely, we reduce the composition of a certain problem (called $\mathsf{pSearch}$ [11]) and the claw problem with the function range stated in Theorem 4 to the claw problem in question, and then use the composition theorem of the general adversary bound [18]. That is, for every $k,F,G\in\mathbb{N}$ with $k|F$ and $G|(F/k)$, the composition $\mathsf{claw}_{(F,G)\to M+2}$ is reducible from the problem $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$, where $\mathsf{claw}_{(a,b)\to c}$ denotes the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Since $\mathsf{pSearch}_{N\to M}$ is an variant of the unstructured search over $N$ elements, the quantum query complexity of $\mathsf{pSearch}_{F/k\to M}$ is $\Omega(\sqrt{F/k})$. By setting $M=k+kG/F$, Theorem 4 provides a lower bound on the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}$ as
$Q_{1/3}\left(\mathsf{claw}_{(k,kG/F)\to M}\right)=\Omega\left(\sqrt{k\frac{G}{F}}+\left(k^{2}\frac{G}{F}\right)^{1/3}\right).$
By applying the compositon theorem of the general adversary bound, we have
$Q_{1/3}\left(\mathsf{claw}_{(F,G)\to M+2}\right)$ $\geq Q_{1/3}\left(\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}\right)$
$=\Omega\left(\left[\sqrt{k\frac{G}{F}}+\left(k^{2}\frac{G}{F}\right)^{1/3}\right]\sqrt{\frac{F}{k}}\right)=\Omega\left(\sqrt{G}+k^{1/6}F^{1/6}G^{1/3}\right).$
We now interpret this in terms of $M$ by substituting $k=M/(G/F+1)=\Theta(MF/G)$ as
$Q_{1/3}\left(\mathsf{claw}_{(F,G)\to M+2}\right)=\Omega\left(\sqrt{G}+\left(\frac{MF}{G}\right)^{1/6}F^{1/6}G^{1/3}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$
Notice that this holds for very special combinations of $k,F,G$ such that $k|F$ and $G|(F/k)$ and $M=k+kG/F$. However, we can remove this constraint and obtain Theorem 5.
1.5 Organization
Sec. missing  2 provides notations and necessary notions. Sec. missing  3 proves Theorem 1 and Theorem 2. Sec. missing  4 proves Theorem 3 and Theorem 4. Sec. missing  5 proves Theorem 5.
2 Preliminaries
2.1 Notations
Let $\mathbb{N}$ be the set of natural numbers, $\mathbb{Z}_{+}$ the set of non-negative integers, $\mathbb{Q}$ the set of rational numbers, $\mathbb{C}$ the set of complex numbers, and $[n]$ the set $\{1,\dots,n\}$ for every $n\in\mathbb{N}$. If two sets $S$ and $T$ are disjoint, we may write their union as $S\sqcup T$ instead of $S\cup T$ to emphasize their disjointness.
For $n\in\mathbb{N}$, let $S_{n}$ be the group of permutations over $[n]$. For a commutative ring $R$ and $n\in\mathbb{N}$, $R[x_{1},\dots,x_{n}]$ denotes the ring of polynomials in $x_{1},\dots,x_{n}$ over $R$. For a polynomial $p\in R[x_{1},\dots,x_{n}]$, we write the degree of $p$ (i.e., the maximum over all product terms in $p$ of the number of variables in a single product term) as $\deg(p)$. We say a polynomial is multilinear if the polynomial is linear in any one of the variables when regarding the other variables as constants. For any $n\times m$ matrix $\Omega=(\omega_{ij})$ over $\mathbb{Z}$, $\lvert\Omega\rvert_{1}$ denotes the sum of the absolute values of all elements in $\Omega$, i.e., $\lvert\Omega\rvert_{1}\equiv\sum_{i=1}^{n}\sum_{j=1}^{m}\lvert\omega_{ij}\rvert$.
[…]
For two functions $(f,g)\in[M]^{[F]}\times[M]^{[G]}$, where $F,G,M\in\mathbb{N}$, we say that a pair $(x,y)\in[F]\times[G]$ is a claw if $f(x)=g(y)$. These definitions are naturally generalized to the case of more than two functions: Let $M\in\mathbb{N}$ and $F_{\ell}\in\mathbb{N}$ for each $\ell\in[k]$, where $k\in\mathbb{N}$ is fixed. Then, for functions $f_{\ell}\in[M]^{[F_{\ell}]}$ for $\ell\in[k]$, we say that $k$-tuple $(x_{1},\dots,x_{k})\in[F_{1}]\times\dots\times[F_{k}]$ is a $k$-claw if $f_{i}(x_{i})=f_{j}(x_{j})$ for all $i,j\in[k]$.
For a function $f$, $\mathrm{dom}(f)$ denotes the domain of $f$, and $f(S)$ denotes the image of a subset $S\subseteq\mathrm{dom}(f)$ via $f$. We can thus write the image of $f$ as $f(\mathrm{dom}(f))$.
Let $\mathcal{D}\subseteq\{0,1\}^{n}$. For a partial Boolean function $\Phi\colon\mathcal{D}\to\{0,1\}$ and a non-negative real $\epsilon\in[0,1/2)$, we say that a polynomial $p\colon\mathbb{R}^{n}\to\mathbb{R}$ $\epsilon$-approximates $\Phi$ if $p$ satisfies $|\Phi(x)-p(x)|\leq\epsilon$ and $p(x)\in[0,1]$ for every $x\in\mathcal{D}$. We say the $\epsilon$-approximate degree of $\Phi$, denotedy by $\widetilde{\deg}_{\epsilon}(\Phi)$, to mean the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$.
The quantum query complexity of a quantum algorithm $\mathcal{A}$ for $\Phi$ with error $\epsilon$, denoted by $Q^{\mathcal{A}}_{\epsilon}(\Phi)$ is the number of quantum queries required by $\mathcal{A}$ to compute $\Phi(x)$ with error probability at most $\epsilon$, when the worst-case input $x\in\mathcal{D}$ is given as oracle. The $\epsilon$-error quantum query complexity of $\Phi$, $Q_{\epsilon}(\Phi)$, is the minimum of $Q^{\mathcal{A}}_{\epsilon}(\Phi)$ over all quantum algorithms $\mathcal{A}$. We assume that readers have basic knowledge of quantum computation.
2.2 Polynomial Method for Quantum Lower bound
The polynomial method was introduced by Beals et al. [9] to obtain lower bounds on the number of quantum queries required to compute a Boolean function on an input bit-string given as an oracle. It is not difficult to generalize the method to the case where the input is a string of alphabets [1]. In the following, we consider the case where the input consists of two alphabet-strings corresponding to two functions. But it is not difficult to see similar facts hold for three or more functions.
The following lemma says that, after $q$ quantum queries, any quantum algorithm is in the state where the amplitude of each computational basis state is represented by a polynomial of degree at most $q$ in $x_{ij}$ and $y_{kj}$.
Lemma 6
Let $F,G,M$ be in $\mathbb{N}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, assume without loss of generality that any quantum algorithm with $q$ queries applies
$U_{q}O_{f/g}U_{q-1}\cdots U_{1}O_{f/g}U_{0}$
to a fixed initial state, say, the all-zero state, where $U_{q},\dots,U_{0}$ are unitary operators independent of the oracles, and $O_{f/g}$ acts as
$|0,i\rangle|b\rangle|z\rangle\rightarrow|0,i\rangle|b+f(i)\rangle|z\rangle,\mbox{ and }|1,i\rangle|b\rangle|z\rangle\rightarrow|1,i\rangle|b+g(i)\rangle|z\rangle,$
where $+$ is the addition modulo $M$, or alternatively,
$|0,i\rangle|b\rangle|z\rangle\rightarrow|0,i\rangle|b\oplus f(i)\rangle|z\rangle,\mbox{ and }|1,i\rangle|b\rangle|z\rangle\rightarrow|1,i\rangle|b\oplus g(i)\rangle|z\rangle,$
where $\oplus$ is bitwise XOR in binary expression. At any step, the the state of the algorithm is of the form
$\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle},$
where the first and second registers consist of $1+\max\{\lceil\log F\rceil,\lceil\log G\rceil\}$ qubits and $\lceil\log{M}\rceil$ qubits, respectively (the last register is a working register that $U_{i}$ acts on but $O_{f/g}$ does not). Define $FM$ variables $x_{ij}\in\{0,1\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$, and $GM$ variables $y_{kj}\in\{0,1\}$ such that $y_{kj}=1$ if and only if $g(k)=j$. Then, in the final state, each $\alpha_{s,i,b,z}$ can be represented as a polynomial over $FM+GM$ variables $x_{ij},y_{kj}$ of degree at most $q$.
The consequence of the polynomial method is as follows.
Lemma 7[9, 1])
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a partial Boolean function over the set of pairs of functions in $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $(f,g)\in\mathcal{D}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec. missing  2.2.
The following lemma says that if $\delta$ and $\epsilon$ are constants, then $\widetilde{\deg}_{\epsilon}(\Phi)$ and $\widetilde{\deg}_{\delta}(\Phi)$ are linearly related.
Lemma 8 ([15])
For a partial Boolean function $\Phi$ and $20<\epsilon<\delta<1/$ 2, it holds that
$\widetilde{\deg}_{\epsilon}(\Phi)\leq\widetilde{\deg}_{\delta}(\Phi)\cdot O\left(\frac{\log(1/\epsilon)}{(1/2-\delta)^{2}}\right).$
In particular, if $\delta$ and $\epsilon$ are constants, then $\widetilde{\deg}_{\epsilon}(\Phi)$ and $\widetilde{\deg}_{\delta}(\Phi)$ are linearly related.
For completeness, the proofs of the above lemmas are given in Appendix A.
2.3 Multisymmetric Properties
For $F,G,M\in\mathbb{N}$, let $\Phi$ be a partial Boolean function over the set of pairs of functions in $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$. Intuitively, we can think of $\Phi$ as an indicator function deciding whether a pair of functions, $(f,g)\in\mathcal{D}$, have a certain property. For instance, we say that $\Phi$ is the claw property, if $\Phi\left(f,g\right)=1$ if and only if there exists a claw for $(f,g)\in[M]^{[F]}\times[M]^{[G]}$. We say that $\Phi$ is bisymmetric, if $\Phi$ is invariant and $\mathcal{D}$ is closed under the permutations over $\mathrm{dom}(f)=[F]$ and $\mathrm{dom}(g)=[G]$, respectively, and the permutation over the common range $[M]$ of $f$ and $g$, i.e., $\left(\sigma f\pi_{F},\sigma g\pi_{G}\right)\in\mathcal{D}$ if $(f,g)\in\mathcal{D}$, and
$\Phi\left(f,g\right)=\Phi\left(\sigma f\pi_{F},\sigma g\pi_{G}\right),$
for all permutations $\pi_{F},\pi_{G},\sigma$ over $[F]$, $[G]$, and $[M]$, respectively. Here $(\sigma f\pi_{F})(i)=\sigma(f(\pi_{F}(i)))$ and $(\sigma g\pi_{G})(j)=\sigma(g(\pi_{G}(j)))$. For instance, if $\Phi$ is the claw property, then $\Phi$ is bisymmetric.
The definition of bisymmetric properties can be generalized to the three or more function case in a straightforward way. Let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a partial Boolean function over the set of $k$-tuples of functions in $\mathcal{D}\subseteq\mathcal{F}_{1}\times\cdots\times\mathcal{F}_{k}$, where $\mathcal{F}_{\ell}\ (\ell\in[k])$ denotes $[M]^{[F_{\ell}]}$ for $F_{\ell},M\in\mathbb{Z}_{+}$. We say that $\Phi$ is $k$-symmetric, if $\Phi(f_{1},\dots,f_{k})$ is invariant and $\mathcal{D}$ is closed under the permutations over $\mathrm{dom}(f_{\ell})=[F_{\ell}]$ for each $\ell\in[k]$, respectively, and the permutation over the common range $[M]$ of $f_{\ell}$’s, i.e., $\left(\sigma f_{1}\pi_{F_{1}},\dots,\sigma f_{k}\pi_{F_{k}}\right)\in\mathcal{D}$ if $(f_{1},\dots,f_{k})\in\mathcal{D}$, and
$\Phi(f_{1},\dots,f_{k})=\Phi\left(\sigma f_{1}\pi_{F_{1}},\dots,\sigma f_{k}\pi_{F_{k}}\right),$
for every permutation $\pi_{F_{\ell}}$ over $[F_{\ell}]$ for $\ell\in[k]$ and every permutation $\sigma$ over $[M]$, respectively.
2.4 Multisymmetric Polynomials
This subsection extends the properties of symmetric polynomials to the case of multisymmetric polynomials.
Recall that a symmetric polynomial over a commutative ring $R$ is a polynomial in $R[x_{1},\dots,x_{n}]$ invariant under the action of any permutation $\sigma\in S_{n}$ on the variables, i.e., the action of ${x}_{i}\mapsto{x}_{\sigma(i)}$. To define multisymmetric polynomials as a generalization of symmetric polynomials, we consider a vector consisting of $m$ variables, $X_{i}:=(x_{i1},\dots,x_{im})$, instead of variable $x_{i}$, for each $i\in[n]$. Then, a multisymmetric polynomial is defined as a polynomial $p\in R[x_{11},\dots,x_{nm}]$ invariant under the action of any permutation $\sigma\in S_{n}$ on vectors $X_{i}$, i.e., the action of $x_{ij}\mapsto x_{\sigma(i)j}$. Let $R[x_{11},\dots,x_{nm}]^{S_{n}}$ be the ring of multisymmetric polynomials in $R[x_{11},\dots,x_{nm}]$. For convenience, let us think of $nm$ variables $x_{11},\dots,x_{nm}$ as the $n\times m$ matrix $X:=(X_{1}^{T},\dots,X_{n}^{T})^{T}$. Then, a mutisymmetric polynomial in $R[x_{11},\dots,x_{nm}]^{S_{n}}$ is invariant under the action of the permutation on rows of $X$.
For ${\Omega}:=({\Omega}_{1}^{T},\dots,{\Omega}_{n}^{T})^{T}\in\mathbb{Z}_{+}^{nm}$ consiting of ${\Omega}_{i}:=(\omega_{i1},\dots,\omega_{im})\in\mathbb{Z}_{+}^{m}$ for each $i\in[n]$, define
$X_{i}^{{\Omega}_{i}}:=\prod_{j\in[m]}x_{ij}^{\omega_{ij}},\ \ \ {X}^{{\Omega}}:=\prod_{i\in[n]}X_{i}^{{\Omega}_{i}}.$
By the definition, ${X}^{{\Omega}}$ is a monomial of degree $\lvert\Omega\rvert_{1}=\sum_{i=1}^{n}\sum_{j=1}^{m}\omega_{ij}$ (note $\lvert\omega_{ij}\rvert=\omega_{ij}$ for $\omega_{ij}\in\mathbb{Z}_{+}$). We also define
$\mathrm{mon}_{\Omega}(X):=\sum_{\Lambda\in S_{n}\Omega}{X}^{\Lambda},$
where $S_{n}\Omega$ is the orbit of $\Omega$ under the action of $\Omega_{i}\mapsto\Omega_{\sigma(i)}$ over all permutations $\sigma\in S_{n}$. For instance, if $X^{\Omega}=x_{12}x_{21}$, then $\mathrm{mon}_{\Omega}(X)=\sum_{i,j\in[n]\colon i\neq j}x_{i2}x_{j1}$.
Observe that the ring $R[x_{11},\dots,x_{nm}]^{S_{n}}$ is a module over $R$ that has a basis consisting of $\mathrm{mon}_{\Omega}(X)$ for ${\Omega}\in\mathbb{Z}_{+}^{nm}$. More concretely, any mutisymmetric polynomial of total degree $d$ in $R[x_{11},\dots,x_{nm}]^{S_{n}}$ is a linear combination of $\mathrm{mon}_{\Omega}(X)$ for $\Omega$’s with $\lvert\Omega\rvert_{1}\leq d$.
Let us define the power-sum $P_{\lambda}(X)$ over $X$ for row vector $\lambda\in\mathbb{Z}_{+}^{m}$ as
$P_{\lambda}(X)=X_{1}^{\lambda}+\dots+X_{n}^{\lambda}.$
The fundamental theorem of multisymmetric polynomials is as follows. The proof is given in Appendix B.
Theorem 9 (e.g. [16])
Let $R$ be a commutative ring that contains $\mathbb{Q}$. For each $\Omega\in\mathbb{Z}_{+}^{nm}$, $\mathrm{mon}_{\Omega}(X)$ can be expressed as a linear combination over $\mathbb{Q}$ of products $P_{\Omega^{\prime}_{1}}(X)\cdots P_{\Omega^{\prime}_{n}}(X)$ of the power-sums over $X=(x_{ij})_{i\in[n],j\in[m]}$ for $\Omega^{\prime}:=((\Omega^{\prime}_{1})^{T},\cdots,(\Omega^{\prime}_{n})^{T})^{T}\in\mathbb{Z}_{+}^{nm}$ such that $\lvert\Omega^{\prime}\rvert_{1}\leq\lvert\Omega\rvert_{1}$. Moreover, the power-sums $P_{\lambda}(X)$ for $\lambda\in\mathbb{Z}_{+}^{m}$ generate the ring $R[x_{11},\dots,x_{nm}]^{S_{n}}$ of multisymmetric polynomials.
3 Polynomials Approximating Multisymmetric Properties
This section proves Theorem 1 by generalizing Ambainis’s proof idea with Theorem 9. Theorem 2 is a straightfoward extension of Theorem 1.
3.1 Bisymmetric Properties
We first show that there are two polynomials in different sets of variables of the same degree such that both polynomials approximates $\Phi$.
Lemma 10
For $F,G,M\in\mathbb{N}$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of functions $(f,g)\in\mathcal{D}$. Define $FM$ variables $x_{ij}$ and $GM$ variables $y_{kj}$ for functions $f$ and $g$ as in Sec. missing  2.2. In addition, define $z_{j}$ as the number of preimages of $j$ via $f$, i.e., $z_{j}=\left|f^{-1}\left(j\right)\right|$, and define $w_{j}$ as the number of preimages of $j$ via $g$, i.e., $w_{j}=\left|g^{-1}\left(j\right)\right|$. Then, the following are equivalent.
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $z_{1},\dots,z_{M},w_{1},\dots,w_{M}$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $P$ of degree at most $d$ in $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ that $\epsilon$-approximates $\Phi$.
Moreover, the minimum degree of a polynomial in $z_{1},\dots,z_{M},w_{1},\dots,w_{M}$ that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial in $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ that $\epsilon$-approximates $\Phi$.
Proof.  The first item implies the second, since $z_{j}=\sum_{i}x_{ij}$ and $w_{j}=\sum_{k}y_{kj}$ by the definitions, and replacing $z_{j}$ with $\sum_{i}x_{ij}$ and $w_{j}$ with $\sum_{k}y_{kj}$ does not change the value and degree of the polynomial.
To show that the second item implies the first, we take the average of $P$, denoted by $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$, over the permutations $\pi_{F}$ of first index $i$ of $x_{ij}$ and the permutations $\pi_{G}$ of the first index $k$ of $y_{kj}$. Since $\Phi$ is invariant under the permutations over the domains of $f,g$, the average $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$ still $\epsilon$-approximates $\Phi$. We will show that the expectation is a polynomial over $z_{j},w_{j}$ of degree at most $d$, assuming that the second item holds.
Since $P$ is a linear combination of terms $x_{i_{1}j_{1}}\dots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\dots y_{k_{l}j_{l}}$ for positive integers $m$ and $l$ with $m+l\leq d$, it suffices to show that the expectation of any single term is a polynomial over $z_{j},w_{j}$ of degree at most $d$. Then, the expectation is
$\mathsf{E}_{\pi_{F}\times\pi_{G}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right]=\mathsf{E}_{\pi_{F}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}\right]\cdot\mathsf{E}_{\pi_{G}}\left[y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right].$
Since we define $\{x_{ij}\}$ as the representations of the function $f$, we do not need to consider invalid assignments to $x_{ij}$’s for which no function can be defined, such as $x_{ij}=x_{ij^{\prime}}=1$ for distinct $j,j^{\prime}$. In addition, since $x_{ij}$ are Boolean, we can assume $x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}$ is multilinear. Therefore, $i_{1},\dots,i_{m}$ are all distinct. Similarly, $k_{1},\dots,k_{\ell}$ are all distinct.
From the analysis in the single function case [4] assuming valid $x_{ij}$ and $y_{kj}$, it holds that $\mathsf{E}_{\pi_{F}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}\right]$ is a polynomial in $z_{j}$ of degree at most $m$ and $\mathsf{E}_{\pi_{G}}\left[y_{i_{1}j_{1}}\cdots y_{i_{l}j_{l}}\right]$ is a polynomial in $w_{j}$ of degree at most $l$ (see Appendix C for the completeness). Therefore, $\mathsf{E}_{\pi_{F}\times\pi_{G}}\left[x_{i_{1}j_{1}}\cdots x_{i_{m}j_{m}}y_{k_{1}j_{1}}\cdots y_{k_{l}j_{l}}\right]$ is a polynomial in $z_{j},w_{j}$ of degree $m+l\leq d$. By defining $\mathsf{E}_{\pi_{F}\times\pi_{G}}[P]$ as $Q$, the second iterm implies the first. This completes the former part of the proof.
For the latter part of the statement, let $d^{\prime}$ be the minimum degree of $P$. By the former part, the degree of $Q$ is at most $d^{\prime}$. If the degree of $Q$ were at most $d^{\prime}-1$, then there would exist a polynomial $P$ of degree at most $d^{\prime}-1$ that $\epsilon$-approximates $\Phi$, which is a contradiction. Thus, the minimum degree of $Q$ is also $d^{\prime}$. $\square$
Next, we show that the degree of a polynomial $Q$ shown in Sec. missing  3.1 does not change, even if we restrict the common range of input functions of $\Phi$ to a small subset.
Lemma 11
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric partial Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$, i.e., the bisymmetric Boolean function over the set of function pairs $(f^{\prime},g^{\prime})\in\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)$ such that $\Phi^{\prime}(f^{\prime},g^{\prime})=\Phi(f,g)$ for all functions $f^{\prime},g^{\prime},f,g$ such that $f^{\prime}(x)=f(x)$ and $g^{\prime}(y)=g(y)$ for every $x\in[F]$ and $y\in[G]$. Define $z_{j}=\left|f^{-1}\left(j\right)\right|$ and $w_{j}=\left|g^{-1}\left(j\right)\right|$ for $f$ and $g$ and define $z^{\prime}_{j}=\left|(f^{\prime})^{-1}\left(j\right)\right|$ and $w^{\prime}_{j}=\left|(g^{\prime})^{-1}\left(j\right)\right|$ for $f^{\prime}$ and $g^{\prime}$. Then, the following are equivalent:
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $z_{1},\dots,z_{M},w_{1},\dots,w_{M}$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $Q^{\prime}$ of degree at most $d$ in $z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$ that $\epsilon$-approximates $\Phi^{\prime}$.
Moreover, the minimum degree of a polynomial in $z_{1},\dots,z_{M},w_{1},\dots,w_{M}$ that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial in $z_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$ that $\epsilon$-approximates $\Phi^{\prime}$.
Proof.  To show that the first item implies the second, let us think of $f^{\prime},g^{\prime}$ as special cases of $(f,g)\in\mathcal{D}$ such that both $f([F])$ and $g([G])$ are contained in $[M^{\prime}]$ (note that such $(f^{\prime},g^{\prime})$ exists in $\mathcal{D}$ since $\mathcal{D}$ is closed under the permutations over the range). In this case, we have $z_{j}=z^{\prime}_{j}$ and $w_{j}=w^{\prime}_{j}$ for all $j\in[M^{\prime}]$, and $z_{i}=w_{j}=0$ for all $j>M^{\prime}$. Thus, it holds that
$\Phi^{\prime}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}})=\Phi(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},0,\dots,0,w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}},0,\dots,0).$
Suppose that the first item holds. If we define
$Q^{\prime}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}):=Q(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},0,\dots,0,w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}},0,\dots,0),$
then $Q^{\prime}$ is a polynomial degree at most $d$ that $\epsilon$-approximates $\Phi^{\prime}$. The second item thus holds.
We next show that the second item implies the first. Fix $(f,g)\in\mathcal{D}$ arbitrarily. Observe then that there exists a permutation $\sigma$ over $[M]$ such that the images of $\sigma f$ and $\sigma g$ are contained in $[M^{\prime}]$, since $\mathrm{dom}(f)$ and $\mathrm{dom}(g)$ have cardinality $F$ and $G$, respectively. Let us define $f^{\prime},g^{\prime}$ by identifying $(f^{\prime},g^{\prime})$ with $(\sigma f,\sigma g)$ as an element in $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, it holds that $z^{\prime}_{j}=z_{\sigma^{-1}(j)}$ and $w^{\prime}_{j}=w_{\sigma^{-1}(j)}$ for every $j\in[M^{\prime}]$. By bisymmetry, we have $\Phi(f,g)=\Phi(\sigma f,\sigma g)$, which is equal to $\Phi^{\prime}(f^{\prime},g^{\prime})$ by the definition.
Suppose that the second item holds. Then, $\Phi^{\prime}(f^{\prime},g^{\prime})$ is $\epsilon$-approximated by some polynomial $Q^{\prime}$ of degree at most $d$ over $z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$. Since $\Phi^{\prime}$ is a bisymmetric property, the average of $Q^{\prime}$ over all permutations over $[M^{\prime}]$, i.e.,
$\widetilde{Q^{\prime}}:=\mathsf{E}_{\tau\in S_{M^{\prime}}}\left[Q^{\prime}(z^{\prime}_{\tau(1)},\dots,z^{\prime}_{\tau(M^{\prime})},w^{\prime}_{\tau(1)},\dots,w^{\prime}_{\tau(M^{\prime})})\right]$
still $\epsilon$-approximates $\Phi^{\prime}$. Since $\widetilde{Q^{\prime}}$ is a bisymmetric polynomial, Theorem 9 implies that $\widetilde{Q^{\prime}}$ is represented as a polynomial, in power-sum polynomials $P^{\prime}_{\lambda}$, of degree at most $d$ with respect to variables $z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}$, where
$P^{\prime}_{\lambda}:=P^{\prime}_{\lambda}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}})=\sum_{i=1}^{M^{\prime}}(z^{\prime}_{i})^{\lambda_{z}}(w^{\prime}_{i})^{\lambda_{w}},$
for $\lambda=(\lambda_{z},\lambda_{w})\in\mathbb{Z}_{+}\times\mathbb{Z}_{+}$ such that $\lambda_{z}+\lambda_{w}\leq d$. Define $P_{\lambda}$ by extending $P^{\prime}_{\lambda}$ to the case $2M$ variables as
$P_{\lambda}:=P_{\lambda}(z_{1},\dots,z_{M},w_{1},\dots,w_{M})=\sum_{i=1}^{M}z_{i}^{\lambda_{z}}w_{i}^{\lambda_{w}}.$
In addition, define a bisymmetric polynomial $Q^{\prime\prime}$ of degree at most $d$ by replacing every occurrence of $P^{\prime}_{\lambda}$ in $\widetilde{Q^{\prime}}$ with $P_{\lambda}$. Then, we have
$Q^{\prime\prime}(z_{1},\dots,z_{M},w_{1},\dots,w_{M})$ $=Q^{\prime\prime}(z_{\sigma^{-1}(1)},\dots,z_{\sigma^{-1}(M)},w_{\sigma^{-1}(1)},\dots,w_{\sigma^{-1}(M)})$
$=Q^{\prime\prime}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},0,\dots,0,w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}},0,\dots,0)$
$=\widetilde{Q^{\prime}}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}),$
where the first equality follows from the bisymmetry of $Q^{\prime\prime}$, the second follows from the definition of permutation $\sigma$, and the last follows from
$P_{\lambda}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},0,\dots,0,w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}},0,\dots,0)=P^{\prime}_{\lambda}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}}).$
This implies that $Q^{\prime\prime}(z_{1},\dots,z_{M},w_{1},\dots,w_{M})$ $\epsilon$-approximates $\Phi(z_{1},\dots,z_{M},w_{1},\dots,w_{M})$, since $\widetilde{Q^{\prime}}$ $\epsilon$-approximates $\Phi^{\prime}(z^{\prime}_{1},\dots,z^{\prime}_{M^{\prime}},w^{\prime}_{1},\dots,w^{\prime}_{M^{\prime}})$, which is equal to $\Phi(z_{1},\dots,z_{M},w_{1},\dots,w_{M})$ by the definition. By setting $Q:=Q^{\prime\prime}$, the first item in the statement holds.
For the latter part of the statement, let $d^{\prime}$ be the minimum degree of $Q$. By the former part, the degree of $Q^{\prime}$ is at most $d^{\prime}$. If the degree of $Q^{\prime}$ were at most $d^{\prime}-1$, then there would exist a polynomial $Q$ of degree at most $d^{\prime}-1$ that $\epsilon$-approximates $\Phi$, which is a contradiction. Thus, the minimum degree of $Q^{\prime}$ is also $d^{\prime}$. $\square$
By combining Sec. missing  3.1 with Sec. missing  3.1, we obtain a slightly general version of Theorem 1.
Theorem 12 (general version of Theorem 1)
For $F,G,M\in\mathbb{N}$ such that $M\geq F+G$ and $\mathcal{D}\subseteq[M]^{[F]}\times[M]^{[G]}$, let $\Phi\colon\mathcal{D}\to\{0,1\}$ be a bisymmetric Boolean function over the set of function pairs $(f,g)\in\mathcal{D}$. In addition, for every $M^{\prime}\in\mathbb{N}$ such that $F+G\leq M^{\prime}<M$, let $\Phi^{\prime}\colon\mathcal{D}\cap\left([M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$. Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$, where
  • $\Phi$ depends on variables $x_{ij}\in\{0,1\}$ for $(i,j)\in[F]\times[M]$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$, and variables $y_{kj}\in\{0,1\}$ for $(k,j)\in[G]\times[M]$ such that $y_{kj}=1$ if and only if $g(k)=j$, and
  • $\Phi^{\prime}$ depends on variables $x^{\prime}_{ij}\in\{0,1\}$ for $(i,j)\in[F]\times[M^{\prime}]$ such that $x^{\prime}_{ij}=1$ if and only if $f^{\prime}\left(i\right)=j$, and variables $y^{\prime}_{kj}\in\{0,1\}$ for $(k,j)\in[G]\times[M^{\prime}]$ such that $y^{\prime}_{kj}=1$ if and only if $g^{\prime}(k)=j$.
With the promise that the union of the images of input functions $f$ and $g$, i.e., ${f([F])\cup g([G])}$, is upper-bounded in size by $\mathfrak{U}$ for every $(f,g)\in\mathcal{D}$, Sec. missing  3.1 holds for every $M^{\prime}$ with $\mathfrak{U}\leq M^{\prime}<M$ with a slight modification to proof. Thus, with that promise, Theorem 12 also holds for such $M^{\prime}$. For instance, if $\mathcal{D}$ is the set of the pairs of two-to-one functions, then $\lvert f([F])\cup g([G])\rvert\leq(F+G)/2$, and thus Theorem 12 holds for every $M^{\prime}$ with $(F+G)/2\leq M^{\prime}<M$. Since $\lvert f([F])\cup g([G])\rvert\leq F+G$ always holds, we obtain Theorem 12 for all $\mathcal{D}$.
3.2 $k$-Symmetric Properties
This section states a straightforward generalization of Lemmas 3.1 and 3.1 to the $k$-symmetric properties. We omit their proofs, since their proofs are almost the same as those in the bisymmetric case. Theorem 15 (a slightly general version of Theorem 2) follows immediately from the following lemmas.
Lemma 13
Let $k\geq 2$. For $F_{\ell}\in\mathbb{N}\ (\ell\in[k])$, $M\in\mathbb{N}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell\in[k]$. For each $\ell\in[k]$, $i\in[F_{\ell}]$, and $j\in[M]$, define variable $x^{(\ell)}_{ij}$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$, and define $z^{(\ell)}_{j}$ as the number of preimages of $j\in[M]$ via $f_{\ell}$, i.e., $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$. Then, the following are equivalent.
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $kM$ variables $z^{(\ell)}_{1},\dots,z^{(\ell)}_{M}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $P$ of degree at most $d$ in $\left(\sum_{\ell=1}^{k}F_{\ell}M\right)$ variables $x^{(\ell)}_{11},\dots,x^{(\ell)}_{F_{\ell}M}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi$.
Moreover, the minimum degree of a polynomial in $z^{(\ell)}_{1},\dots,z^{(\ell)}_{M}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial in $x^{(\ell)}_{11},\dots,x^{(\ell)}_{F_{\ell}M}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi$.
Lemma 14
. Let $k\geq 2$. For $F_{\ell}\ (\ell\in[k])$ and $M\in\mathbb{N}$ such that $M\geq F_{1}+\dots+F_{k}$, let $\Phi:\mathcal{D}\to\{0,1\}$ be a $k$-symmetric partial Boolean function over the set of $k$-tuples of functions $(f_{1},\dots,f_{k})\in\mathcal{D}$, where $\mathcal{D}\subseteq\mathcal{F}_{1}\times\dots\times\mathcal{F}_{k}$ and $\mathcal{F}_{\ell}$ denotes $[M]^{[F_{\ell}]}$ for each $\ell\in[k]$. In addition, for every ${M^{\prime}}\in\mathbb{N}$ such that $F_{1}+\dots+F_{k}\leq{M^{\prime}}<M$, let $\Phi^{\prime}:\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)\to\{0,1\}$ be the restriction of $\Phi$ to the domain $[{M^{\prime}}]^{[F_{1}]}\times\dots\times[{M^{\prime}}]^{[F_{k}]}$, i.e., the $k$-symmetric Boolean function over the set of $k$-tuples of functions $(f^{\prime}_{1},\dots,f^{\prime}_{k})\in\mathcal{D}\cap\left(\mathcal{F}^{\prime}_{1}\times\dots\times\mathcal{F}^{\prime}_{k}\right)$, where $\mathcal{F}^{\prime}_{\ell}$ denotes $[{M^{\prime}}]^{[F_{\ell}]}$ for each $\ell\in[k]$, such that $\Phi^{\prime}(f^{\prime}_{1},\dots,f^{\prime}_{k})=\Phi(f_{1},\dots,f_{k})$ for all functions $f^{\prime}_{1},\dots,f^{\prime}_{k},f_{1},\dots,f_{k}$ satisfying $f^{\prime}_{\ell}(x)=f_{\ell}(x)$ for all $x\in[F_{\ell}]$ and all $\ell\in[k]$. Finally, for each $\ell\in[k]$, define $z^{(\ell)}_{j}=\left|f_{\ell}^{-1}\left(j\right)\right|$ for each $j\in[M]$ and ${z^{\prime}}^{(\ell)}_{j}=\left|(f^{\prime}_{\ell})^{-1}\left(j\right)\right|$ for each $j\in[{M^{\prime}}]$. Then, the following are equivalent:
  1. 1.
    There exists a polynomial $Q$ of degree at most $d$ in $kM$ variables $z^{(\ell)}_{1},\dots,z^{(\ell)}_{M}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi$.
  2. 2.
    There exists a polynomial $Q^{\prime}$ of degree at most $d$ in ${z^{\prime}}^{(\ell)}_{1},\dots,{z^{\prime}}^{(\ell)}_{{M^{\prime}}}\ (\ell\in[k])$ that $\epsilon$-approximates $\Phi^{\prime}$.
Theorem 15 (general version of Theorem 2)
Define $\Phi$ and $\Phi^{\prime}$ as in Sec. missing  3.2, where
  • $\Phi$ depends on $x^{(\ell)}_{ij}\in\{0,1\}$ for $(\ell,i,j)\in[k]\times[F_{\ell}]\times[M]$ such that $x^{(\ell)}_{ij}=1$ if and only if $f_{\ell}\left(i\right)=j$,
  • $\Phi^{\prime}$ depends on ${x^{\prime}}^{(\ell)}_{ij}\in\{0,1\}$ for $(\ell,i,j)\in[k]\times[F_{\ell}]\times[{M^{\prime}}]$ such that ${x^{\prime}}^{(\ell)}_{ij}=1$ if and only if ${f}^{\prime}_{\ell}\left(i\right)=j$.
Then, the minimum degree of a polynomial that $\epsilon$-approximates $\Phi$ is equal to the minimum degree of a polynomial that $\epsilon$-approximates $\Phi^{\prime}$.
4 Degree of Claw Property
This section provides a lower bound for the claw problem with a small range using Theorem 1. For this, we need a lower bound on the minimum degree of a polynomial that approximates the claw property for some common range size of input functions. To prove the following theorem, we use the Zhang’s reduction [32] from the collision problem for an input function that is either one-to-one or two-to-one.
Theorem 16
For any constant $20<\epsilon<1/$ 2, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $[FG]^{F}\times[FG]^{G}$ is $\Omega\left((FG)^{1/3}\right)$.
Proof.  The basic idea is as follows. Suppose that there exists a polynomial $P^{\prime}$ of degree $d$ that 0.1-approximates the claw property for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$ such that $F\leq G$. Based on $P^{\prime}$, we will then construct a polynomial $P$ of degree $d$ that (41/86)-approximates the collision property for functions $h\in[M]^{[M]}$ for $M=FG$ with the promise that $h$ is either one-to-one or two-to-one. Here we mean by the collision property the Boolean function $\Phi$ which is $\mathsf{true}$ if and only if the input function $h$ is two-to-one. Since the minimum degree of the latter polynomial is $\Omega(M^{1/3})$ [2, 4], we obtain $d=\Omega(M^{1/3})$. Below, we make this idea more precise.
Take a pair of disjoint subsets $S\subset[M]$ and $T\subset[M]$ of size $F$ and $G$, respectively, and consider the restrictions of $h$ to $S$ and $T$, $h|_{S}\in[M]^{S}$ and $h|_{T}\in[M]^{T}$, respectively. Here, we assume that $h$ depends on Boolean variables $x_{ij}$ for $i,j\in[M]$ such that $x_{ij}=1$ if and only if $h(i)=j$. Accordingly, $h|_{S}$ depends on $x_{ij}$ for $i\in S$ and $j\in[M]$; $h|_{T}$ depends on $x_{ij}$ for $i\in T$ and $j\in[M]$. Let $p_{\rm cl}(h)$ be the probability that $h|_{S}$ and $h|_{T}$ have a claw over a pair of disjoint subsets $S$ and $T$ (of fixed size $F$ and $G$, respectively) chosen uniformly at random.
For disjoint subsets $S,T\subset[M]$, let $P^{\prime}_{ST}$ be a polynomial of degree $d$ (in variables $x_{ij}$ for $i\in S\sqcup T$ and $j\in[M]$) that $\epsilon$-approximates the claw property on function pairs $[M]^{S}\times[M]^{T}$. For every function $h$, let $P^{\prime}_{ST}(h)$ denote the value of $P^{\prime}_{ST}$ at the assignment to $x_{ij}$ corresponding to $h|_{S}$ and $h|_{T}$. Then, define a polynomial $P$ as the average of $P^{\prime}_{ST}$ over uniformly random disjoint subsets $S,T\subset[M]$ of size $F$ and $G$, respectively:
$P:=\mathsf{E}_{S,T\subset[M]\colon S\cap T=\emptyset}[P^{\prime}_{ST}].$
Thus, $P$ is a polynomial of degree at most $d$ in variables $x_{ij}$ for $i,j\in[M]$. Let $P(h)$ denote the average of $P^{\prime}_{ST}(h)$ over $S,T$ for every function $h$.
If $h$ is one-to-one, then $h|_{S}$ and $h|_{T}$ obviously have no claws for every disjoint $S$ and $T$, i.e., $p_{\rm cl}(h)=0$. Thus, it holds $0\leq P^{\prime}_{ST}(h)\leq\epsilon$ for all disjoint $S,T\subset[M]$. This implies
$0\leq P(h)\leq\epsilon.$
If $h$ is two-to-one, then $P^{\prime}_{ST}(h)\geq 1-\epsilon$ for a $p_{\rm cl}(h)$ fraction of all pairs of disjoint subsets $S$ and $T$. This implies
$P(h)\geq p_{\rm cl}(h)(1-\epsilon).$
If we set $M=FG$ for $G\geq F\geq 400$ and $\epsilon=0.1$, then it follows from Sec. missing  4 that $P(h)>0.2\cdot(1-0.1)=0.18$. Therefore, $P$ separates the one-to-one and two-to-one cases with a constant gap. By scaling and shifting $P$, we have a polynomial that $O(1)$-approximates the collision property. For instance, $\tilde{P}:=(25P+18)/43$ is sufficient and its approximation error is at most $41/86$. Therefore, $\deg(\tilde{P})=\Omega(M^{1/3})=\Omega((FG)^{1/3})$. Since $d\geq\deg(P)=\deg(\tilde{P})$, the minimum degree $d$ of a polynomial that 0.1-approximates the claw property for function pairs $[FG]^{F}\times[FG]^{G}$ is $\Omega\left((FG)^{1/3}\right)$. By Sec. missing  2.2, the theorem holds. $\square$
Claim 17
Suppose that $h\in[M]^{[M]}$ is two-to-one. For $M=FG$ and $G\geq F\geq 400$, it holds that $p_{\rm cl}(h)>0.2$.
Proof.  To choose random disjoint subsets $S$ and $T$, we first choose a random subset $S$ of $[M]$ and then choose a random subset $T$ of $[M]\setminus S$. Observe that, if $h|_{S}$ is injective and, in addition, $h^{-1}(h(S))\setminus S$ intersects $T$, then a claw exists. Thus, we obtain
$p_{\rm cl}(h)>\Pr_{S,T\subset[M]\colon S\cap T=\emptyset}\left[(h|_{S}\text{ is injective})\wedge(T\text{ intersects }h^{-1}(h(S))\setminus S)\right].$
We will derive a bound on $M$ that satisfies
$\Pr_{S\subset[M]}\left[h|_{S}\text{ is injective}\right]\geq 1-\delta\ \text{and}\ \Pr_{T\subset[M]\setminus S}\left[(T\text{ intersects }h^{-1}(h(S))\setminus S)\mid(h|_{S}\text{ is injective})\right]\geq 1-\delta^{\prime},$
so that
$p_{\rm cl}(h)>(1-\delta)(1-\delta^{\prime})>1-(\delta+\delta^{\prime}).$
The probability $\Pr_{S}\left[h|_{S}\text{ is injective}\right]$ is
$\frac{M(M-2)\cdots(M-2(F-1))}{F!\binom{M}{F}}$ $=\frac{M(M-2)\cdots(M-2(F-1))}{M(M-1)\cdots(M-F+1)}$
$=1\cdot\left(1-\frac{1}{M-1}\right)\left(1-\frac{2}{M-2}\right)\cdots\left(1-\frac{F-1}{M-F+1}\right)$
$>\prod_{j=1}^{F-1}\exp\left(-\frac{j}{M-2j}\right)$
$>\exp\left(-\frac{F(F-1)}{2}\frac{1}{M-2F}\right)>\exp\left(-\frac{F^{2}}{2M-F^{2}/100}\right),$
where the first inequality follows from $1+x\geq\exp(\frac{x}{1+x})$, and the last inequality follows from $1004F\leq F^{2}/$ 100 for $F\geq 400$. For the last formula to be at least $1-\delta$, we have
$M$ $\geq\frac{F^{2}}{2}\left[\frac{1}{\ln\left(\frac{1}{1-\delta}\right)}+\frac{1}{100}\right].$ (1)
Next, we consider a bound on $M$ such that
$\Pr_{T\subset[M]\setminus S}\left[(T\text{ intersects }h^{-1}(h(S))\setminus S)\mid(h|_{S}\text{ is injective})\right]=1-\frac{\binom{M-2F}{G}}{\binom{M-F}{G}}$
is lower-bounded by $1-\delta^{\prime}$. Equivalently, we will derive a bound on $M$ such that ${\binom{M-2F}{G}}/{\binom{M-F}{G}}\leq\delta^{\prime}$.
$\frac{\binom{M-2F}{G}}{\binom{M-F}{G}}$ $=\frac{(M-2F)(M-2F-1)\cdots(M-2F-(G-1))}{(M-F)(M-F-1)\cdots(M-F-(G-1))}$
$=\left(1-\frac{F}{M-F}\right)\left(1-\frac{F}{M-F-1}\right)\cdots\left(1-\frac{F}{M-F-(G-1)}\right)$
$<\prod_{j=0}^{G-1}\exp\left(-\frac{F}{M-F-j}\right)<\exp\left(-\frac{FG}{M}\right).$
For this to be at most $\delta^{\prime}$, we have
$M\leq\frac{FG}{\ln({1}/{\delta^{\prime}})}.$ (2)
From Ineqs. (1) and (2), if there exist $M,\delta,\delta^{\prime}$ such that $\delta+\delta^{\prime}<1$ and
$\frac{F^{2}}{2}\left[\frac{1}{\ln\left(\frac{1}{1-\delta}\right)}+\frac{1}{100}\right]\leq\frac{FG}{2}\left[\frac{1}{\ln\left(\frac{1}{1-\delta}\right)}+\frac{1}{100}\right]\leq M\leq\frac{FG}{\ln({1}/{\delta^{\prime}})},$ (3)
then $p_{\rm cl}(h)>1-(\delta+\delta^{\prime})>0.$
For $\delta=\delta^{\prime}=0.4$, Ineq. (3) is $0.98\ldots\times FG\leq M\leq 1.09...\times FG$. Therefore, it holds $p_{\rm cl}(h)>1-(0.4+0.4)=0.2$ for $M=FG$. $\square$
As a corollary of Theorem 12, we obtain the following theorem.
Theorem 18 (restatement of Theorem 3)
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For any constant $20<\epsilon<1/$ 2 and any $M\geq F+G$, the minimum degree of a polynomial that $\epsilon$-approximates the claw property for function pairs in $[M]^{[F]}\times[M]^{[G]}$ is $\Theta(\sqrt{G}+(FG)^{1/3})$, where the lower bound $\Omega(\sqrt{G})$ holds for all $M\geq 2$.
Proof.  Let $\Phi$ be the claw property for functions in $[FG]^{[F]}\times[FG]^{[G]}$. By Theorem 16, we have $\widetilde{\deg}_{\epsilon}(\Phi)=\Omega((FG)^{1/3})$ when $M=FG$. By applying Theorem 12, we have $\widetilde{\deg}_{\epsilon}(\Phi^{\prime})=\Omega((FG)^{1/3})$, where $\Phi^{\prime}$ is the restriction of $\Phi$ to the domain $[M^{\prime}]^{[F]}\times[M^{\prime}]^{[G]}$ for any $M^{\prime}$ with $F+G\leq M^{\prime}<FG$. This proves $\Omega((FG)^{1/3})$ lower bound in the statement of the theorem.
The $\Omega(\sqrt{G})$ lower bound follows using the reduction from the unstructured search [14]: Let $p$ be a polynomial of $\deg_{\epsilon}(\Phi)$ in variables $x_{ij},y_{kj}$ for $(i,k,j)\in[F]\times[G]\times[M]$ that $\epsilon$-approximates the claw property $\Phi$ on a function pair $(f,g)\in[M]^{[F]}\times[M]^{[G]}$. Suppose $M\geq 2$. Then consider the special case where $f$ maps all inputs to $1$ and $g$ maps all inputs to either $1$ or $2$. This is equivalent to setting $x_{i1}=1$ for all $i\in[F]$, $x_{ij}=0$ for all $i,j$ with $j\geq 2$, and $y_{kj}=0$ for all $k,j$ with $j\geq 3$. Then, the polynomial $p^{\prime}$ resulting from $p$ with these replacements depends only on $y_{kj}$ over $(k,j)\in[G]\times[2]$. Since we can assume $y_{k2}=1-y_{k1}$ in this special case, $p^{\prime}$ can be rewritten as a polynomial of the same degree in $y_{k1}$ over $k\in[G]$. Observe now that, if there exists $k$ such that $y_{k1}=1$ (meaning the existence of a claw for $f,g$), it holds $1-\epsilon\leq p^{\prime}\leq 1$; otherwise, there exists no claw and thus $0\leq p^{\prime}\leq\epsilon$. This implies that $p^{\prime}$ $\epsilon$-approximates the OR function over $G$ variables $y_{k1}$’s. Because of $\widetilde{\deg}_{\epsilon}(OR_{G})=\Theta(\sqrt{G})$ [28], we have $\widetilde{\deg}_{\epsilon}(\Phi)\geq\deg{p^{\prime}}=\Omega(\sqrt{G})$.
Since there is an $O\left(\sqrt{G}+(FG)^{1/3}\right)$-query quantum algorithm that detects a claw (and finds it if it exists) [30], the polynomial method (Sec. missing  2.2) implies that the lower bound $\Omega\left(\sqrt{G}+(FG)^{1/3}\right)$ is tight for $M\geq F+G$. $\square$
The proof of Theorem 18 implies the optimal quantum query complexity as follows.
Theorem 19 (restatement of Theorem 4)
Let $F,G\in\mathbb{N}$ be such that $F\leq G$. For every $M\geq F+G$, the (query-)optimal quantum algorithm that computes a claw (or detects the existence of a claw) with a constant error probability requires $\Theta(\sqrt{G}+(FG)^{1/3})$ queries for a given pair of functions in $[M]^{[F]}\times[M]^{[G]}$. More precisely, the optimal quantum query complexity is $\Theta((FG)^{1/3})$ for every $M\geq F+G$ if $F\leq G\leq F^{2}$ and $\Theta(\sqrt{G})$ for every $M\geq 2$ if $G>F^{2}$.
5 Claw Problem with Much Smaller Range
This section proves Theorem 5. The proof generalizes the idea used in the case of the same domain size [7]: We first reduce the composition of a certain problem (called $\mathsf{pSearch}$) and the claw problem for a larger size of the function range to the claw problem in question, and then use the composition theorem of the adversary method.
Theorem 20 ([18])
For finite sets $C$ and $E$, and $\mathcal{D}\subseteq C^{n}$, let $f\colon\mathcal{D}\to E$. An adversary matrix for $f$ is a $\lvert\mathcal{D}\rvert\times\lvert\mathcal{D}\rvert$ real, symmetric matrix $\Gamma$ that satisfies $\langle x|\Gamma|y\rangle=0$ for all $x,y\in\mathcal{D}$ with $f(x)=f(y)$. Define the general adversary bound as
$\mathsf{ADV}^{\pm}(f):=\max_{\Gamma}\frac{\|\Gamma\|}{\max_{j\in[n]}\|\Gamma\circ\Delta_{j}\|},$
where the maximization is over all possible adversary matrices $\Gamma$ of $f$, and where $\Gamma\circ\Delta_{j}$ denotes the entry-wise matrix product between $\Gamma$ and $\Delta_{j}=\sum_{x,y\in D\colon x_{j}\neq y_{j}}|x\rangle\langle y|$. Then, the two-sided $\epsilon$-bounded error quantum query complexity of $f$ is lower bounded by $\mathsf{ADV}^{\pm}(f)$ as follows:
$Q_{\epsilon}(f)\geq\frac{1-2\sqrt{\epsilon(1-\epsilon)}}{2}\mathsf{ADV}^{\pm}(f).$
Definition 21 ($\mathsf{pSearch}$ [11])
$\mathsf{pSearch}_{K\to M}$ denotes the problem defined as follows. For $K,M\in\mathbb{N}$, given an oracle in $([M]\cup\{\ast\})^{K}$ with the promise that exactly one of the $K$ values is non-$\ast$, the goal is to output the non-$\ast$ value.
Since $\mathsf{pSearch}_{K\to M}$ can be computed with $O(\sqrt{K})$ queries by using Grover search and it requires $\Omega(\sqrt{K})$ quantum queries (by the adversary method [3]),
$Q_{1/3}(\mathsf{pSearch}_{K\to M})=\Theta\left(\sqrt{K}\right).$ (4)
Brassard et al. [11] proved the following composition lemma.
Lemma 22 ([11, Theorem 9])
For any function $f\colon[M]^{N}\to Z$ and $\mathsf{pSearch}_{K\to M}$, define $h\colon D\to Z$, where $D\subset([M]\cup\{\ast\})^{KN}$ such that $((x_{11},\dots,x_{1K}),\dots,(x_{N1},\dots,x_{NK}))$ is in $D$ if and only if exactly one of $x_{i1},\dots,x_{iK}$ is non-$\ast$ for every $i\in[N]$, as $h:=f\circ\mathsf{pSearch}_{K\to M},$ that is,
$h\left((x_{11},\dots,x_{1K}),\dots,(x_{N1},\dots,x_{NK})\right):=f\left(\mathsf{pSearch}_{K\to M}(x_{11},\dots,x_{1K}),\dots,\mathsf{pSearch}_{K\to M}(x_{N1},\dots,x_{NK})\right).$
Then,
$\mathsf{ADV}^{\pm}(h)\geq\frac{2}{\pi}\mathsf{ADV}^{\pm}(f)\cdot\mathsf{ADV}^{\pm}(\mathsf{pSearch}_{K\to M}).$
By employing Theorem 19 and Sec. missing  5, we obtain the following theorem.
Theorem 23 (restatement of Theorem 5)
Let $F,G,M\in\mathbb{N}$ be such that $F\leq G\leq F^{2}$ and $M<F+G$ [^2] The condition of $F\leq G\leq F^{2}$ and $M<F+G$ means the parameter range that Theorem 19 does not cover. Then, for every $M\in[2,F+G-1]$, the quantum query complexity of detecting the existence of a claw for a given function pair $(f,g)$ in $[M]^{[F]}\times[M]^{[G]}$ is lower-bounded by
$\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)=\left\{\begin{array}[]{lcl}\Omega(\sqrt{G})&\text{if}&2\leq M<(G/F)^{2},\\ \Omega(F^{1/3}G^{1/6}M^{1/6})&\text{if}&(G/F)^{2}\leq M<F+G.\end{array}\right.$
Proof.  We will show that the following problem $\mathcal{P}$ is reducible to the claw problem and then give a lower bound on the quantum query complexity of $\mathcal{P}$.
The input to $\mathcal{P}$ is a pair of functions $f:=(f_{1},\dots,f_{F})\in([M]\cup\{\ast\})^{[F]}$ and $g:=(g_{1},\dots,g_{G})\in([M]\cup\{\ast\})^{[G]}$ (as the sequence representation of functions; see Sec. missing  2.1) for $M\geq 2$ with the following promise, where $f_{i}$ denotes $f(i)$ for $i\in[F]$ and $g_{k}$ denotes $g(k)$ for $k\in[G]$:
  • if $f$ is partitioned into $k$ blocks of size $F/k$,
    $(f_{1},\dots,f_{F/k}),\dots,(f_{F-F/k+1},\dots,f_{F}),$
    then each block contains exactly one non-$\ast$ value;
  • if $g$ is partitioned into $kG/F$ blocks of size $F/k$,
    $(g_{1},\dots,g_{F/k}),\dots,(g_{G-F/k+1},\dots,g_{G}),$
    then each block contains exactly one non-$\ast$ value.
Here, we assume the promise that $k$ and $F/k$ divide $F$ and $G$, respectively. This assumption will be removed later.
The goal of $\mathcal{P}$ is to accept if, when applying $\mathsf{pSearch}$ on each block of $f$ and $g$, the collection of the outputs of $\mathsf{pSearch}$ for $f$ intersects the collection of the outputs of $\mathsf{pSearch}$ for $g$, and reject otherwise. Equivalently, the goal is to detect the existence of a claw for two functions identified with the following two sequences, respectively:
$\left(\mathsf{pSearch}_{F/k\to M}(f_{1},\dots,f_{F/k}),\dots,\mathsf{pSearch}_{F/k\to M}(f_{F-F/k+1},\dots,f_{F})\right),$
$\left(\mathsf{pSearch}_{F/k\to M}(g_{1},\dots,g_{F/k}),\dots,\mathsf{pSearch}_{F/k\to M}(g_{G-F/k+1},\dots,g_{G})\right).$
Observe that the problem $\mathcal{P}$ for $f,g$ with $k|F$ and $(k/F)|G$ is reducible to detecting the existence of a claw for function $f^{\prime}\in[M+2]^{[F]}$ and $g^{\prime}\in[M+2]^{[G]}$, where $f^{\prime}$ and $g^{\prime}$ are obtained by replacing every occurrence of $\ast$ with $M+1$ and $M+2$, respectively. That is, for each $(f,g)$ with the promise and the corresponding $(f^{\prime},g^{\prime})$, it holds that
$\mathsf{claw}_{(F,G)\to M+2}(f^{\prime},g^{\prime})=\left(\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}\right)(f,g),$
where $\mathsf{claw}_{(a,b)\to c}$is the decision problem of detecting the existence of a claw for given functions $(\phi,\psi)\in[c]^{[a]}\times[c]^{[b]}$. Note that, since each query access to $f^{\prime}\ (g^{\prime})$ can be simulated by a single query access to $f\ (\text{respectively},\ g)$, the modification to input functions needs no extra queries. Therefore, the quantum query complexity of $\mathsf{claw}_{(F,G)\to M+2}$ is lower-bounded by the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}$.
Let us set $M=k+kG/F$, for which we will bound the quantum query complexity of $\mathsf{claw}_{(k,kG/F)\to M}$ by using Theorem 19: If $k^{2}<kG/F$ (i.e., $k<G/F$), then $Q(\mathsf{claw}_{(k,kG/F)\to M})=\Omega(\sqrt{kG/F})$. If $kG/F\leq k^{2}$ (i.e., $G/F\leq k$), then $Q(\mathsf{claw}_{(k,kG/F)\to M})=\Omega\left((k\cdot kG/F)^{1/3}\right)$. More succinctly,
$Q_{1/3}\left(\mathsf{claw}_{(k,kG/F)\to M}\right)=\Omega\left(\sqrt{k\frac{G}{F}}+\left(k^{2}\frac{G}{F}\right)^{1/3}\right).$ (5)
By applying Theorem 20 and Sec. missing  5 with Eq. (missing) 4 and Eq. (missing) 5, we have
$Q_{1/3}\left(\mathsf{claw}_{(F,G)\to M+2}\right)$ $\geq Q_{1/3}\left(\mathsf{claw}_{(k,kG/F)\to M}\circ\mathsf{pSearch}_{F/k\to M}\right)$
$=\Omega\left(\left[\sqrt{k\frac{G}{F}}+\left(k^{2}\frac{G}{F}\right)^{1/3}\right]\sqrt{\frac{F}{k}}\right)=\Omega\left(\sqrt{G}+k^{1/6}F^{1/6}G^{1/3}\right).$
We now interpret this in terms of $M$ by substituting $k=M/(G/F+1)=\Theta(MF/G)$ as
$Q_{1/3}\left(\mathsf{claw}_{(F,G)\to M+2}\right)=\Omega\left(\sqrt{G}+\left(\frac{MF}{G}\right)^{1/6}F^{1/6}G^{1/3}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right),$ (6)
where $M=k+kG/F$.
To remove the assumption that $k$ and $F/k$ divides $F$ and $G$, respectively, let us consider any $F,G,k$ such that $F\leq G\leq F^{2}$ and $1\leq k\leq F$. Define
$F^{\prime}:=k\left\lfloor\frac{F}{k}\right\rfloor,\ \ \ \ \ \ \ G^{\prime}:=\frac{F^{\prime}}{k}\left\lfloor\frac{G}{F^{\prime}/k}\right\rfloor.$
Note that
$G^{\prime}=\frac{F^{\prime}}{k}\left\lfloor k\frac{G}{F^{\prime}}\right\rfloor\geq\frac{F^{\prime}}{k}k=F^{\prime}.$
By the definition, $k$ divides $F^{\prime}$ and $F^{\prime}/k$ divides $G^{\prime}$. Then, for any instance $f^{\prime}\colon[F^{\prime}]\to[M+2]$ and $g^{\prime}\colon[G^{\prime}]\to[M+2]$ of $\mathsf{claw}_{(F^{\prime},G^{\prime})\to M+2}$, we define $f\colon[F]\to[M+2]$ and $g\colon[G]\to[M+2]$ as
  • $f(x):=f^{\prime}(x)$ for all $x\in[F^{\prime}]$, and $f(x):=f^{\prime}(F^{\prime})$ for all $x\in[F]\setminus[F^{\prime}]$;
  • $g(y):=g^{\prime}(y)$ for all $y\in[G^{\prime}]$, and $g(y):=g^{\prime}(G^{\prime})$ for all $y\in[G]\setminus[G^{\prime}]$.
Then, one can see that there exists a claw for $\left(f^{\prime},g^{\prime}\right)$ if and only if there exists a claw for $\left(f,g\right)$. For $M=k+kG^{\prime}/F^{\prime}$, we thus have
$Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M+2}\right)$ $\geq Q_{1/3}\left(\mathsf{claw}_{\left(F^{\prime},G^{\prime}\right)\rightarrow M+2}\right)$
$=\Omega\left(\sqrt{G}+\left(F^{\prime}\right)^{1/3}({G^{\prime})\ }^{1/6}M^{1/6}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right),$
where the first equality follows from Eq. (missing) 6 and the last equality follows from $F^{\prime}\geq F/2$ and $G^{\prime}\geq G/2$ as shown in Sec. missing  5.
Next, we will remove the assumption that $M$ is of the form $M=k+kG^{\prime}/F^{\prime}$ and show that the lower bound holds for every integer $M$ such that $2\leq M<F+G$. For this, let us fix $F$ and $G$ arbitrarily such that $F\leq G\leq F^{2}$. For every integer $M$ such that $M\in\left[1+\left\lfloor G/F\right\rfloor,F+G\right]$, there exists $k\in[1,F]$ such that $M_{k}\leq M\leq M_{k+1}$ by Sec. missing  5, where $M_{k}$ is defined in the claim. Since $\mathsf{claw}_{\left(F,G\right)\rightarrow M_{k}+2}$ is reducible to $\mathsf{claw}_{\left(F,G\right)\rightarrow M+2}$ since $[M_{k}+2]\subseteq[M+2]$, we have
$Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M+2}\right)\geq Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M_{k}+2}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M_{k}^{1/6}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right),$
where the last equality follows from $M/M_{k}<M_{k+1}/M_{k}\leq 8$ by Sec. missing  5. By replacing $M+2$ with $M$, we have
$Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$
for $3+\left\lfloor G/F\right\rfloor\leq M\leq F+G$. Since $\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)=\Omega\left(\sqrt{G}\right)$ for every positive $M<3+\left\lfloor G/F\right\rfloor$ and $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}\right)$ [14] for every $M\geq 2$ follows from the unstructured search, it holds that $Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\Omega\left(\sqrt{G}+F^{1/3}G^{1/6}M^{1/6}\right)$ for $2\leq M\leq F+G$. More precisely, Since $\sqrt{G}>F^{1/3}G^{1/6}M^{1/6}$ if and only if $\left(G/F\right)^{2}>M$, we have [^3] The case $2\leq M<(G/F)^{2}$ should be ignored when $G/F\leq\sqrt{2}$.
$Q_{1/3}\left(\mathsf{claw}_{\left(F,G\right)\rightarrow M}\right)=\left\{\begin{array}[]{lcl}\Omega\left(\sqrt{G}\right)&\text{if}&2\leq M<\left(G/F\right)^{2}\\ \Omega\left(F^{1/3}G^{1/6}M^{1/6}\right)&\text{if}&\left(G/F\right)^{2}\leq M<F+G.\end{array}\right.$
$\square$
Claim 24
For every $F,G\in\mathbb{N}$ such that $F\leq G\leq F^{2}$, let us define $M_{k}:=k+kG_{k}^{\prime}/F_{k}^{\prime}$ for every $1\leq k\leq F$, where $F_{k}^{\prime}=k\left\lfloor F/k\right\rfloor$ and $G_{k}^{\prime}=(F_{k}^{\prime}/k\ )\left\lfloor\frac{G}{\left(F_{k}^{\prime}/k\right)}\right\rfloor$. Then, it holds that
  1. 1.
    $F/2\leq F_{k}^{\prime}\leq F$, and $G/2\leq G_{k}^{\prime}\leq G$ for every $k\in\left[1,F\right]$,
  2. 2.
    $M_{k}\leq M_{k+1}$ for every $k\in\left[1,F-1\right]$,
  3. 3.
    $M_{1}=1+\left\lfloor G/F\right\rfloor$ and $M_{F}=F+G$,
  4. 4.
    $M_{k+1}/M_{k}\leq 8$ for every $k\in\left[1,F-1\right]$.
Proof.  We first prove the second and third items. Since $kG_{k}^{\prime}/F_{k}^{\prime}=\left\lfloor\frac{G}{\left(F^{\prime}/k\right)}\right\rfloor$ is the number of blocks of size $F_{k}^{\prime}/k$ in $G^{\prime}$ and the block size $F_{k}^{\prime}/k=\left\lfloor F/k\right\rfloor$ is monotone decreasing with respect to $k$, $kG_{k}^{\prime}/F_{k}^{\prime}$ is monotone increasing. Therefore, $M_{k}$ is monotone increasing with respect to $k$. Therefore, $M_{k}\leq M_{k+1}$ for every $k\in\left[1,F-1\right]$. Since $F_{1}^{\prime}=F$ and $G_{1}^{\prime}=\left\lfloor G/F\right\rfloor$, we have $M_{1}=1+\left\lfloor G/F\right\rfloor$. Since $F_{F}^{\prime}=F$ and $G_{F}^{\prime}=G$, we have $M_{F}=F+G$.
We next prove the first and the last items. Since $F_{k}^{\prime}\leq F$ and $G_{k}^{\prime}\leq G$ by the definition, the first item follows by proving the lower bounds on $F^{\prime}_{k}$ and $G^{\prime}_{k}$. By the definition of $F_{k}^{\prime}$, we have $F-F_{k}^{\prime}<\left\lfloor F/k\right\rfloor$ and $F_{k}^{\prime}\geq\left\lfloor F/k\right\rfloor$. Hence,
$F_{k}^{\prime}\geq\max{\left\{F-\left\lfloor\frac{F}{k}\right\rfloor,\left\lfloor\frac{F}{k}\right\rfloor\right\}}\geq\frac{F}{2}.$
By the definition of $G_{k}^{\prime}$, we have $G-G_{k}^{\prime}<F_{k}^{\prime}/k=\left\lfloor F/k\right\rfloor$ and $G_{k}^{\prime}\geq F_{k}^{\prime}/k=\left\lfloor F/k\right\rfloor$. Hence,
$G_{k}^{\prime}\geq\max{\left\{G-\left\lfloor\frac{F}{k}\right\rfloor,\left\lfloor\frac{F}{k}\right\rfloor\right\}}\geq\frac{G}{2}.$
Together with $F_{k}^{\prime}\leq F$ and $G_{k}^{\prime}\leq G$, we have
$\frac{1}{2}\cdot\frac{G}{F}\leq\frac{G_{k}^{\prime}}{F_{k}^{\prime}}\leq 2\frac{G}{F}.$
Then, the last item follows from
$\frac{M_{k+1}}{M_{k}}=\frac{k+1}{k}\frac{1+G_{k+1}^{\prime}/F_{k+1}^{\prime}}{1+G_{k}^{\prime}/F_{k}^{\prime}}\leq 2\cdot\frac{1+2G/F}{1+G/\left(2F\right)}\leq 8.$
$\square$
Acknowledgments
This work was partially supported by JSPS KAKENHI Grant Numbers JP20H05966 and JP22H00522.
Appendix
Appendix A Claim of Polynomial Method
Lemma 25 (restatement of Sec. missing  2.2)
Let $F,G,M$ be in $\mathbb{N}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, assume without loss of generality that any quantum algorithm with $q$ queries applies
$U_{q}O_{f/g}U_{q-1}\cdots U_{1}O_{f/g}U_{0}$
to the initial state, say, the all-zero state, where $U_{q},\dots,U_{0}$ are unitary operators independent of the oracles, and $O_{f/g}$ acts as
$|0,i\rangle|b\rangle|z\rangle\rightarrow|0,i\rangle|b+f(i)\rangle|z\rangle,\mbox{ and }|1,i\rangle|b\rangle|z\rangle\rightarrow|1,i\rangle|b+g(i)\rangle|z\rangle,$
where $+$ is the addition modulo $M$, or alternatively,
$|0,i\rangle|b\rangle|z\rangle\rightarrow|0,i\rangle|b\oplus f(i)\rangle|z\rangle,\mbox{ and }|1,i\rangle|b\rangle|z\rangle\rightarrow|1,i\rangle|b\oplus g(i)\rangle|z\rangle,$
where $\oplus$ is bitwise XOR in binary expression. At any step, the the state of the algorithm is of the form
$\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle},$
where the first and second registers consist of $1+\max\{\lceil\log F\rceil,\lceil\log G\rceil\}$ qubits and $\lceil\log{M}\rceil$ qubits, respectively (The last register is a working register that $U$ acts on but $O_{f/g}$ does not). Define $FM$ variables $x_{ij}\in\{0,1\}$ such that $x_{ij}=1$ if and only if $f\left(i\right)=j$, and $GM$ variables $y_{kj}\in\{0,1\}$ such that $y_{kj}=1$ if and only if $g(k)=j$. Then, in the final state, each $\alpha_{s,i,b,z}$ can be represented as a polynomial over $FM+GM$ variables $x_{ij},y_{kj}$ of degree at most $q$, where $q$ is the number of quantum queries.
Proof.  For $q=0$, since $\alpha_{s,i,b,z}$ is independent of the oracles, the statement holds. Suppose that the claim holds for $q-1$ and let $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ be the state after applying $U_{q-1}$, where each $\alpha_{s,i,b,z}$ is a polynomial over $x_{ij},y_{ij}$ of degree at most $q-1$. Assume that the $q$th $O_{f/g}$ maps this state to $\sum_{s,i,b,z}{\beta_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$. Observe that $O_{f/g}$ maps each basis state $|0,i\rangle|b\rangle|z\rangle$ to $|0,i\rangle|b+f(i)\rangle|z\rangle$ (likewise $|1,k\rangle|b\rangle|z\rangle$) to $|0,k\rangle|b+g(k)\rangle|z\rangle$), and that $|0,i\rangle|b-j+f(i)\rangle|z\rangle$ is exactly $|0,i\rangle|b\rangle|z\rangle$ if $x_{ij}=1$ (and likewise $|1,k\rangle|b-j+g(k)\rangle|z\rangle$ is exactly $|1,k\rangle|b\rangle|z\rangle$ if $y_{kj}=1$). Thus, it holds that
$\beta_{0,i,b,z}$ $=x_{i1}\alpha_{0,i,b-1,z}+x_{i2}\alpha_{0,i,b-2,z}+\dots+x_{iM}\alpha_{0,i,b-M,z},$
$\beta_{1,k,b,z}$ $=y_{k1}\alpha_{1,k,b-1,z}+y_{k2}\alpha_{1,k,b-2,z}+\dots+y_{kM}\alpha_{1,k,b-M,z}.$
Therefore, $\beta_{0,i,b,z}$ and $\beta_{1,k,b,z}$ are polynomials over $x_{ij}$ and $y_{ij}$ of degree at most $q$. After applying unitary $U_{q}$, the amplitude of each basis state is a linear combination of $\beta_{0,i,b,z}$ and $\beta_{1,k,b,z}$. Therefore, its degree is still at most $q$. In the case of the XOR-type oracle, the same argument holds. $\square$
Lemma 26 (restatement of Sec. missing  2.2)
Let $F,G,M$ be in $\mathbb{N}$, and let $\Phi$ be a Boolean function over the set of pairs of functions in $[M]^{[F]}\times[M]^{[G]}$. For given oracles $O_{f}$ and $O_{g}$ for functions $f\in[M]^{[F]}$ and $g\in[M]^{[G]}$, suppose that there exists a quantum algorithm that computes $\Phi(f,g)$ with error $\epsilon$ using $q$ queries. Then, there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$, where $x_{ij}$ and $y_{kj}$ are the variables defined for $f$ and $g$, respectively, in Sec. missing  2.2.
Proof.  Assume without loss of generality that the quantum algorithm measures the last single-qubit register in the final state $\sum_{s,i,b,z}{\alpha_{s,i,b,z}\left|s,i\right\rangle\left|b\right\rangle\left|z\right\rangle}$ in the computational basis and outputs the outcome as the value of $\Phi$. Let $A$ be the set of tuples $(s,i,b,z)$ with the last bit of $z$ being $\Phi(f,g)$. Then, the probability that the algorithm outputs $\Phi(f,g)$ is
$p_{\Phi}(f,g):=\sum_{(s,i,b,z)\in A}\left\lvert\alpha_{s,i,b,z}\right\rvert^{2}.$
It follows from Sec. missing  2.2 that $p_{\Phi}$ is represented as a polynomial of degree at most $2q$ in $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$. Since the error probability of the algorithm is at most $\epsilon$, we have $\left\lvert p_{\Phi}(f,g)-\Phi(f,g)\right\rvert\leq\epsilon$ for all $f,g$. This implies that there is a polynomial $P$ of degree at most $2q$ in variables $x_{11},\dots,x_{FM},y_{11},\dots,y_{GM}$ such that $P$ $\epsilon$-approximates $\Phi$. $\square$
Lemma 27 ([15]; restatement of Sec. missing  2.2 )
For a partial Boolean function $\Phi$ and $20<\epsilon<\delta<1/$ 2, it holds that
$\widetilde{\deg}_{\epsilon}(\Phi)\leq\widetilde{\deg}_{\delta}(\Phi)\cdot O\left(\frac{\log(1/\epsilon)}{(1/2-\delta)^{2}}\right).$
In particular, if $\delta$ and $\epsilon$ are constants, then $\widetilde{\deg}_{\epsilon}(\Phi)$ and $\widetilde{\deg}_{\delta}(\Phi)$ are linearly related.
Proof.  Let $p$ be a polynomial of $\widetilde{\deg}_{\delta}(\Phi)$ that $\delta$-approximate $\Phi$. Define a degree-$(\ell\cdot\deg_{\delta}(\Phi))$ polynomial $A(p)$ as
$A(p):=\sum_{k\geq\ell/2}\binom{\ell}{k}p^{k}(1-p)^{\ell-k}.$
Intuitively, this is the probability of seeing at least $\ell/2$ heads among $\ell$ coin flips, if $p$ is the probability of seeing the heads when flipping a single coin. By Chernoff’s bound, the approximation error of $A(p)$ is $\exp(-2\ell(1/2-\delta)^{2})$. For this to be $\epsilon$, we have
$\ell=O\left(\frac{\log(1/\epsilon)}{(1/2-\delta)^{2}}\right).$
The latter part follows from the trivial fact that $\widetilde{\deg}_{\delta}(\Phi)\leq\widetilde{\deg}_{\epsilon}(\Phi)$. $\square$
Appendix B Fundamental Theorem of Multisymmetric Polynomials
Theorem 28 (e.g. [16]; restatement of Theorem 9)
Let $R$ be a commutative ring that contains $\mathbb{Q}$. For each $\Omega\in\mathbb{Z}_{+}^{nm}$, $\mathrm{mon}_{\Omega}(X)$ can be expressed as a linear combination over $\mathbb{Q}$ of products of the power-sums, $P_{\Omega^{\prime}_{1}}(X)\cdots P_{\Omega^{\prime}_{n}}(X)$ over $X=(x_{ij})_{i\in[n],j\in[m]}$, for $\Omega^{\prime}:=((\Omega^{\prime}_{1})^{T},\cdots,(\Omega^{\prime}_{n})^{T})^{T}\in\mathbb{Z}_{+}^{nm}$ such that $\lvert\Omega^{\prime}\rvert_{1}\leq\lvert\Omega\rvert_{1}$. Moreover, the power-sums $P_{\lambda}(X)$ for $\lambda\in\mathbb{Z}_{+}^{m}$ generate the ring $R[x_{11},\dots,x_{nm}]^{S_{n}}$ of multisymmetric polynomials.
Proof.  Since every mutisymmetric polynomial of total degree $d$ in $R[x_{11},\dots,x_{nm}]^{S_{n}}$ is generated by $\mathrm{mon}_{\Omega}(X)$ for $\Omega$’s with $\lvert\Omega\rvert_{1}\leq d$, the former part implies that every multisymmetric polynomial can be expressed as a polynomial in the power-sums $P_{\lambda}(X)$ over $R$. Conversely, every polynomial generated by power-sums $P_{\lambda}(X)$ is obviously in $R[x_{11},\dots,x_{nm}]^{S_{n}}$. Therefore, the latter part of the statement holds. It thus suffices to prove the former part.
For each $\Omega=(\Omega_{1}^{T},\dots,\Omega_{n}^{T})^{T}$, consider the product of power-sums:
$P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)=(X_{1}^{\Omega_{1}}+\cdots+X_{n}^{\Omega_{1}})\cdots(X_{1}^{\Omega_{n}}+\cdots+X_{n}^{\Omega_{n}}).$ (7)
If $\Omega=O$, then $P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)=n^{n}$ and $\mathrm{mon}_{\Omega}(X)=1$. Thus, it holds that $\mathrm{mon}_{\Omega}(X)=(1/n^{n})P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)$.
Suppose $\Omega\neq O$. In this case, there is at least one non-zero row. We claim that $P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)$ is represented as
$P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)=c_{\Omega}\cdot\mathrm{mon}_{\Omega}(X)+\sum_{\Omega^{\prime}}c_{\Omega^{\prime}}\cdot\mathrm{mon}_{\Omega^{\prime}}(X),$ (8)
where $c_{\Omega}$ and $c_{\Omega^{\prime}}$ are positive integers (possibly depending on $n$), and the latter summation is taken over $\Omega^{\prime}=((\Omega^{\prime}_{1})^{T},\dots,(\Omega^{\prime}_{n})^{T})^{T}$ such that (1) the number of non-zero rows in $\Omega^{\prime}$ is strictly less than the number of those in $\Omega$, (2) $\lvert\Omega\rvert_{1}=\lvert\Omega^{\prime}\rvert$.
With Eq. (missing) 8, a simple induction on the number of non-zero rows of $\Omega$ completes the proof: Suppose that $\Omega$ has a single non-zero rows, say $\Omega_{i}$. In this case,
$P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)=n^{n-1}(X_{1}^{\Omega_{i}}+\dots+X_{n}^{\Omega_{i}})=n^{n-1}\cdot\mathrm{mon}_{\Omega}(X).$
For $\Omega$ with $k\geq 2$ non-zero rows, it follows from the claim that
$\mathrm{mon}_{\Omega}(X)=\frac{1}{c_{\Omega}}P_{\Omega_{1}}(X)\cdots P_{\Omega_{n}}(X)-\sum_{\Omega^{\prime}}\frac{c_{\Omega^{\prime}}}{c_{\Omega}}\cdot\mathrm{mon}_{\Omega^{\prime}}(X).$
By the induction hypothesis, this implies that $\mathrm{mon}_{\Omega}(X)$ is a linear combination of $P_{\Omega^{\prime}_{1}}(X)\cdots P_{\Omega^{\prime}_{n}}(X)$ with rational coefficients over $\Omega^{\prime}=((\Omega^{\prime}_{1})^{T},\dots,(\Omega^{\prime}_{n})^{T})^{T}$ such that $\lvert\Omega^{\prime}\rvert_{1}\leq\lvert\Omega\rvert_{1}$. This proves the statement in the lemma.
We now prove Eq. (missing) 8. First, consider the sum of the products obtained by taking a term for distinct row vector $X_{i}$ from each clause in the R.H.S. of Eq. (missing) 7. The sum is represented as
$\sum_{\sigma\in S_{n}}X_{1}^{\Omega_{\sigma(1)}}\dots X_{n}^{\Omega_{\sigma(n)}}.$
By the orbit-stabilizer theorem, each distinct term in this sum appears exactly $\lvert(S_{n})_{\Omega}\rvert$ times, where $(S_{n})_{\Omega}$ is the subgroup of $S_{n}$ that stabilizes $\Omega$. Thus, the sum is represented as
$\left\lvert(S_{n})_{\Omega}\right\rvert\sum_{\Lambda\in S_{n}\Omega}X_{1}^{{\Lambda_{1}}}\dots X_{n}^{\Lambda_{n}}=\left\lvert(S_{n})_{\Omega}\right\rvert\cdot\mathrm{mon}_{\Omega}(X).$
Next, consider the remaining products, i.e., the products obtained by taking terms for the same row of $X$ from multiple clause in the L.H.S. of Eq. (missing) 7. For instance, consider the product obtained by taking $X_{i}^{\Omega_{1}}$ and $X_{i}^{\Omega_{2}}$ from the first and second clauses, respectively, and a term for distinct row vector $X_{j}\ (i\neq j)$ from each of the other clause. Then, the product is of the form of $X_{i}^{\Omega_{1}}X_{i}^{\Omega_{2}}Q$, where $Q$ is the product of remaining variables. This can be regarded as $X^{\Omega^{\prime}}$ with $\Omega^{\prime}_{i}=\Omega_{1}+\Omega_{2}$. In general, one can observe that such a product is represented as $X^{\Omega^{\prime}}$, where each non-zero row in $\Omega^{\prime}$ is the sum of a disjoint set of non-zero rows in $\Omega$ so that the number of non-zero rows in $\Omega^{\prime}$ is strictly less than the number of those in $\Omega$, and $\lvert\Omega\rvert_{1}=\lvert\Omega^{\prime}\rvert$. It is not difficult to see that the set of $\Omega^{\prime}$ induced by products obtained by expanding R.H.S. of Eq. (missing) 7 is partitioned into orbits under the action of $S_{n}$ such that $\Omega^{\prime}_{i}\to\Omega^{\prime}_{\sigma(i)}$ for each $i\in[n]$ and $\sigma\in S_{n}$. For any representative $\Omega^{\prime}$ of each orbit, the sum of the products in Eq. (missing) 7 associated with the orbit is expressed by the orbit-stabilizer theorem as
$\left\lvert(S_{n})_{\Omega^{\prime}}\right\rvert\sum_{\Lambda\in S_{n}\Omega^{\prime}}X_{1}^{{\Lambda_{1}}}\dots X_{n}^{\Lambda_{n}}=\left\lvert(S_{n})_{\Omega^{\prime}}\right\rvert\cdot\mathrm{mon}_{\Omega^{\prime}}(X).$
$\square$
Appendix C Symmetrization
For any function $f\in[M]^{[F]}$, the expectation of $x_{\pi(i_{1})j_{1}}\cdots x_{\pi(i_{m})j_{m}}$ over all permutations $\pi$ is
$\mathsf{E}_{\pi}\left[x_{\pi(i_{1})j_{1}}\cdots x_{\pi(i_{m})j_{m}}\right]$ $=1\cdot\Pr_{\pi}{\left[x_{\pi(i_{1})j_{1}}\cdots x_{\pi(i_{m})j_{m}}=1\right]}$
$=\Pr_{\pi}{\left[x_{\pi(i_{1})j_{1}}=1\right]}\cdot\prod_{\ell=2}^{m}\Pr_{\pi}{\left[x_{\pi(i_{\ell})j_{\ell}}=1\mid x_{\pi(i_{1})j_{1}}\cdots x_{\pi(i_{\ell-1})j_{\ell-1\ }}=1\right]},$
where the expectation and the probability are taken over the permutation $\pi$ over $[F]$. Then, $\Pr_{\pi}{\left[x_{\pi(i_{1})j_{1}}=1\right]}=z_{j_{1}}/F$ follows from the definition of $z_{j_{1}}$. Next, we impose on $\pi$ the condition that $f\left(\pi(i_{1})\right)=j_{1},\ldots,f\left(\pi(i_{\ell-1})\right)=j_{\ell-1}$, i.e., $x_{i_{1}j_{1}}\cdots x_{i_{\ell-1}j_{\ell-1\ }}=1$. Then, there exists $F-\left(\ell-1\right)$ choices for $\pi(i_{\ell})$, since the term does not contain the pair of variable with the same first index. To satisfy $x_{\pi(i_{\ell})j_{\ell}}=1$, it must hold that $\pi(i_{\ell})\in f^{-1}\left(j_{\ell}\right)\setminus\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$. Let $s_{\ell}$ be the number of indices in $\{\pi(i_{1}),\ldots,\pi(i_{\ell-1})\}$ such that they are mapped via $f$ to $j_{\ell}$. Then, there are $z_{j_{\ell}}-s_{\ell}$ choices for $\pi(i_{\ell})$ satisfying $x_{\pi(i_{\ell})j_{\ell}}=1$. Therefore,
$\Pr{\left[x_{i_{\ell}j_{\ell}}=1\mid x_{i_{1}j_{1}}\cdots x_{i_{\ell-1}j_{\ell-1\ }}=1\right]}=\frac{z_{j_{\ell}}-s_{\ell}}{F-\left(\ell-1\right)}.$
Hence, $\mathsf{E}_{\pi}\left[x_{\pi(i_{1})j_{1}}\cdots x_{\pi(i_{m})j_{m}}\right]=\prod_{\ell=1}^{m}\frac{z_{j_{\ell}}-s_{\ell}}{F-\left(\ell-1\right)}$, which is a polynomial in $z_{j}$’s of degree at most $m$.